### v1

# Large Language Model Evaluation via Matrix Nuclear-Norm

Abstract
[…]
1 Introduction
Large language models (LLMs), such as Gemini (Gemini et al., 2023), Llama (Touvron et al., 2023), and GPT-4 (GPT-4 Achiam et al., 2023), have demonstrated remarkable performance across a variety of natural language processing (NLP) tasks (Zhao et al., 2023). They are not only transforming the landscape of NLP (Saul et al., 2005; Liu et al., 2023; Sawada et al., 2023) but also bring beneficial impacts on computer vision (Lian et al., 2023a; Wang et al., 2024) and graph neural networks (Zhang et al., 2024; Chen et al., 2024b), achieving stellar results on various leaderboards. Despite these advancements, assessing a model’s ability to compress information remains a crucial research challenge (Delétang et al., 2023). Compression focuses on efficiently distilling essential information from vast training datasets while discarding redundant elements, showcasing a model’s ability to learn and recognize the underlying structure of the data (Wei et al., 2024). LLMs are expected to perform this form of compression during their training process (Zhao et al., 2023). Specifically, in the early stages of training, after random initialization, the representations produced from the data are often chaotic. However, as training progresses, these representations become more organized, allowing the model to filter out unnecessary information. Hence, assessing an LLM’s capacity for information compression is crucial for understanding its learning efficiency and representational power.
Current compression evaluation methods, such as Matrix Entropy introduced by Wei et al. (2024), measure information compression efficiency through processing models’ output representations on datasets. However, the reliance of Matrix Entropy on Singular Value Decomposition (SVD) (Kung et al., 1983; Zhang, 2015) leads to significant computational complexity, typically $O(n^{3})$, which limits its applicability in large-scale models.
[…]
To validate the effectiveness of the Matrix Nuclear-Norm, we first conducted preliminary experiments on two language models of differing sizes. The results indicated a consistent decrease in Matrix Nuclear-Norm values as model size increased, signifying enhanced compression capabilities. Subsequently, we performed inference experiments on two widely used benchmark datasets, AlpacaEval (Dubois et al., 2024) and Chatbot Arena (Chiang et al., 2024), which cover a diverse range of language generation tasks. These benchmarks facilitate a comprehensive assessment of model inference performance. Our experimental findings confirm that the Matrix Nuclear-Norm accurately measures model compression capabilities and effectively ranks models based on performance, demonstrating its reliability and efficiency in practical applications. Our empirical investigations yield the following insights:
  1. 1.
    Proposal of the Matrix Nuclear-Norm: We present a new method that leverages the nuclear norm, successfully reducing the computational complexity associated with evaluating language models from $O(n^{3})$ to $O(n^{2})$. This reduction minimizes dependence on SVD, making the Matrix Nuclear-Norm a more efficient alternative to Matrix Entropy.
  2. 2.
    Extensive Experimental Validation: We validated the effectiveness of the Matrix Nuclear-Norm on language models of various sizes. Results indicate that this metric accurately assesses model compression capabilities, with values decreasing as model size increases, reflecting its robust evaluation capability.
  3. 3.
    Benchmark Testing and Ranking: We conducted inference tests on widely used benchmark datasets, AlpacaEval and Chatbot Arena, evaluating the inference performance of models across different sizes and ranking them based on the Matrix Nuclear-Norm. The results demonstrate that this metric can efficiently and accurately evaluate the inference performance of medium and small-scale models, highlighting its broad application potential in model performance assessment.
2 Related Work
Evaluating Large Language Models: Metrics and Innovations. The evaluation of large language models (LLMs) is rapidly evolving, driven by diverse tasks, datasets, and benchmarks (Wang et al., 2023; Celikyilmaz et al., 2020; Zheng et al., 2023; Chang et al., 2024). Effective evaluation is crucial for enhancing model performance and reliability. Traditional metrics, such as accuracy, F1 score (Sasaki, 2007), BLEU (Papineni et al., 2002), and ROUGE (Lin, 2004), assess model outputs against annotated labels, while perplexity and cross-entropy loss focus on input data. Recent innovations, like Matrix Entropy (Wei et al., 2024), provide a means to evaluate information compression through activation matrix entropy, shedding light on internal representations. However, the computational intensity of Matrix Entropy, reliant on SVD, highlights the urgent requirement for developing more efficient evaluation methods in the context of LLMs.
Scaling Laws and Performance Metrics. Recent studies have consistently shown that as models and datasets expand, performance metrics such as loss improve predictably (Hestness et al., 2017; Kaplan et al., 2020; Clark et al., 2022). This trend is observed across various domains, including NLP and computer vision (Hoffmann et al., 2022; Bahri et al., 2024). Consequently, both industry and academia have prioritized the development of larger models (Wei et al., 2022; Touvron et al., 2023). Notably, scaling laws reveal a power-law relationship in performance enhancement. In addition to traditional metrics like loss and task accuracy (Zhang et al., 2022; Dey et al., 2023; Biderman et al., 2023), we introduce the Matrix Nuclear-Norm as a novel metric for assessing information compression and redundancy reduction in LLMs.
Insights from Information Theory for Large Language Models. Information theory has significantly enhanced our understanding of neural networks. The information bottleneck concept has been pivotal in elucidating supervised learning dynamics (Tishby et al., 2000). Recently, researchers have applied information-theoretic principles to self-supervised learning in vision tasks (Tan et al., 2023; Skean et al., 2023). Additionally, lossless compression techniques for LLMs, utilizing arithmetic coding, have been explored (Chen et al., 2024a; Valmeekam et al., 2023; Delétang et al., 2023). In this paper, we build on these foundations to introduce the Matrix Nuclear-Norm, an information-theoretic metric aimed at effectively evaluating the information compression and redundancy reduction capabilities of LLMs.
3 Preliminaries
This section presents the fundamental concepts and metrics used in our study to assess model performance, specifically focusing on discriminability, diversity, and the nuclear norm.
3.1 Discriminability Measurement: F-norm
Higher discriminability corresponds to lower prediction uncertainty in the response matrix $A$, quantified using Shannon Entropy (Shannon, 1948):
$H(A)=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{C}A_{i,j}\log(A_{i,j}).$ (1)
In this equation, $B$ represents the number of samples, while $C$ denotes the number of distinct categories. Minimizing $H(A)$ results in maximal discriminability (Vu et al., 2019; Long et al., 2016), achieved when each row contains only one entry equal to 1 and all other entries equal to 0, indicating complete prediction certainty. Alternatively, we can focus on maximizing the Frobenius norm $\|A\|_{F}$:
$\|A\|_{F}=\sqrt{\sum_{i=1}^{B}\sum_{j=1}^{C}|A_{i,j}|^{2}}.$ (2)
The Frobenius norm serves as a measure of the overall magnitude of the entries in $A$ and also reflects prediction certainty.
Theorem 1. Given a matrix $A$, $H(A)$ and $\|A\|_{F}$ are strictly inversely monotonic.
The proof is detailed in Sect A.4 of the Appendix. Thus, minimizing $H(A)$ is equivalent to maximizing $\|A\|_{F}$. The bounds for $\|A\|_{F}$ are:
$\sqrt{\frac{B}{C}}\leq\|A\|_{F}\leq\sqrt{B},$ (3)
where the upper and lower bounds correspond to maximum and minimum discriminability, respectively.
3.2 Diversity Measurement: Matrix Rank
Prediction diversity involves the number of unique categories in matrix $A$, denoted as $C_{p}(A)$. The expected value $E_{C}$ represents the average number of categories:
$E_{C}=E_{A\sim\mathcal{D}}\left(C_{p}(A)\right).$ (4)
In practice, $C_{p}(A)$ can be approximated by the rank of $A$ when $\|A\|_{F}$ is near its upper bound:
$C_{p}(A)\approx\operatorname{rank}(A).$ (5)
3.3 Nuclear Norm
The nuclear norm is an important measure related to diversity and discriminability.
Theorem 2. When $\|A\|_{F}\leq 1$, the convex envelope of $\operatorname{rank}(A)$ is the nuclear norm $\|A\|_{\star}$. The theorem is proved in Fazel (2002).
The nuclear norm $\|A\|_{\star}$, defined as the sum of singular values of $A$, has significant implications for assessing model performance. With $\|A\|_{F}\leq\sqrt{B}$, we have:
$\frac{1}{\sqrt{D}}\|A\|_{\star}\leq\|A\|_{F}\leq\|A\|_{\star}\leq\sqrt{D}\cdot\|A\|_{F},$ (6)
where $D=\min(B,C)$. Therefore, maximizing $\|A\|_{\star}$ ensures high diversity and discriminability.
The upper bound of $\|A\|_{\star}$ is given by:
$\|A\|_{\star}\leq\sqrt{D\cdot B}.$ (7)
4 Methodology
4.1 Motivation
[…]
4.2 Matrix Nuclear-Norm
Calculating the nuclear norm of matrix $A$ requires calculating its SVD, with a time complexity of $O\left(\min(B^{2}C,BC^{2})\right)$, simplifying to $O(n^{3})$, where $n$ represents the maximum of dimensions $B$ and $C$. While this computation is manageable for smaller dimensions, it becomes increasingly time-consuming as dimensions grow. Furthermore, SVD may not converge in some cases, it is thus important to develop more efficient approximations of singular values.
Since $A$ is typically sparse, with only a few non-zero responses in each category, its singular values can be approximated by aggregating these non-zero responses.
Theorem 3. When $\|A\|_{F}$ approaches the upper bound $\sqrt{B}$, the $j$-th largest singular value $\sigma_{j}$ can be approximated as:
$\sigma_{j}\approx\text{top}\left(\sum_{i=1}^{B}A_{i,j}^{2},j\right),\quad\forall j\in\{1,\ldots,D\}.$ (8)
The proof is provided in Sect. A.5 of the Supplementary Materials. Using this approximation, the batch nuclear norm can be efficiently calculated as:
$\|\hat{A}\|_{*}=\sum_{j=1}^{D}\operatorname{top}\left(\sqrt{\sum_{i=1}^{B}A_{i,j}^{2}},j\right).$ (9)
This approach indicates that the primary components of the $L_{1,2}$-norm can effectively approximate the nuclear norm when $\|A\|_{F}$ is close to $\sqrt{B}$, while other components can be considered noise. Compared to traditional SVD-based methods, such as Guo et al. (2015), our Matrix Nuclear-Norm calculation offers two main advantages: (1) it reduces computational complexity from $O(n^{3})$ to $O(n^{2})$, and (2) it relies solely on ordinary floating-point calculations, avoiding risks of non-convergence. The complete algorithm for Matrix Nuclear-Norm is detailed in Algorithm 1.
Definition of Matrix Nuclear-Norm. The approach can ultimately be expressed as:
[…]
Here, $L_{\text{input}}$ denotes the length of the input sequence, ensuring comparability through normalization. Our observations indicate that Matrix Nuclear-Norm values increase with longer sequences; further details can be found in Section 5.3.2.
Algorithm 1 Algorithm of Matrix Nuclear-Norm
0:  Sentence representations (hidden states from LLM) $\mathcal{S}=\left\{X_{i}\right\}_{i=1}^{m}$, where $x_{i}\in\mathbb{R}^{d\times 1}$, $d$ is the hidden dimension of representation, and $L_{\text{input}}$ is the length of the sentence.
1:  $\mu=\frac{1}{m}\sum_{i=1}^{m}x_{i}$ // Calculate the mean embedding
2:  $\mathbf{X}_{\text{norm}}=\frac{\mathbf{X_{i}}-\mu}{\|\mathbf{X_{i}}-\mu\|_{2,\text{row}}}$ // Normalize the activation matrix
[…]
6:  return Matrix Nuclear-Norm
5 Experiments of Large Language Models
5.1 Implementation Details
5.1.1 Baselines
Cross-Entropy Loss. For a text sequence $U=\{u_{1},\ldots,u_{T}\}$, the cross-entropy loss is defined as follows (Yang, 2019):
$L(U)=-\frac{1}{T}\sum_{i=1}^{T}\log P(u_{i}|u_{<i};\Theta)$ (11)
This loss quantifies the model’s ability to predict $u_{i}$ based on preceding words $u_{<i}$. We evaluate the performance of the Matrix Entropy metric in comparison to this baseline, utilizing the same datasets and models as referenced in (Kaplan et al., 2020).
Perplexity. The perplexity of a language model for a text sequence $U=\{u_{1},\ldots,u_{T}\}$ is defined as (Neubig, 2017):
$PPL(U)=\exp\left(-\frac{1}{T}\sum_{i=1}^{T}\log P(u_{i}|u_{<i};\Theta)\right)$ (12)
Perplexity measures the model’s capacity to predict the next word, reflecting the average number of attempts required for correct predictions.
Matrix Entropy of a Dataset. Given a dataset $\mathcal{D}=\{\mathcal{S}_{i}\}_{i=1}^{n}$, where $\mathcal{S}_{i}$ represents the samples (sentence embeddings) within the dataset, the (normalized) matrix entropy of $\mathcal{D}$ is defined as (Wei et al., 2024):
$H(\mathcal{D})=\frac{\sum_{i=1}^{n}H(\Sigma_{\mathcal{S}_{i}})}{n\log d}.$ (13)
5.1.2 Language Models
In our experiments, we selected a range of widely used transformer-based LLMs. Notably, we included Cerebras-GPT (Gao et al., 2020), a pre-trained model well-suited for studying scaling laws. The selection of Cerebras-GPT is particularly advantageous due to its diverse model sizes, which span from 111 million to 13 billion parameters. This diversity allows for a comprehensive analysis of pre-trained language models across varying scales. Additionally, we utilized various scaled versions of the Pythia model (Biderman et al., 2023), ranging from 14 million to 12 billion parameters, to further examine performance variations as model scale changes, thus validating the effectiveness of the proposed Matrix Nuclear-Norm metric.
We conducted Matrix Nuclear-Norm calculations and comparative analyses on inference responses from these models using two benchmark datasets: AlpacaEval and ChatBot Arena. The specific models included in our study are the DeepSeek series (Guo et al., 2024) (1.3B, 6.7B, 7B), the Llama3 series (Dubey et al., 2024) (8B, 70B), the QWEN 2 series (Yang et al., 2024) (0.5B, 1.5B, 7B, 72B), and the Vicuna series (Chiang et al., 2023) (7B, 13B, 33B). We also evaluated models of the same scale, specifically Gemma-7B (Team et al., 2024) and Mistral-7B (Jiang et al., 2023). The inclusion of these diverse models enriches our research perspective and facilitates an in-depth exploration of the inference performance and scaling laws of LLMs across different parameter sizes.
5.1.3 Language Datasets
Pretraining Datasets. All sizes of the Cerebras-GPT models were pretrained on the Pile dataset. Due to resource constraints, we randomly selected a subset of 10,000 data samples from Wikipedia (Foundation, 2024) and openwebtext2 (Skylion007, 2019) to observe the Matrix Nuclear-Norm, as these datasets are part of the Pile. In addition to the datasets used for model pretraining, we incorporated supplementary datasets that were not directly involved in the training process.
Instruction-Tuning Datasets. To assess the effectiveness of the Matrix Nuclear-Norm, we selected the dolly-15k dataset (Conover et al., 2023), generated by human workers, as one of our instruction datasets. We specifically focused on the "context" portion of this dataset, which contains more informative text. This dataset will also be referenced in Section 5.3.2 for analyzing context length, prompt learning, and contextual operations.
RLHF Dataset. We utilized the hh-rlhf dataset (Bai et al., 2022), which contains human preference data related to usefulness and harmlessness, as our Reinforcement Learning with Human Feedback (RLHF) dataset. Each entry includes a pair of texts, one marked as "chosen" and the other as "rejected." We input the "chosen" portions of the dataset into the model and subsequently calculated the performance of the Matrix Nuclear-Norm.
Benchmark Datasets. To evaluate the performance of the Matrix Nuclear-Norm, we selected three benchmark datasets: OpenBookQA (Mihaylov et al., 2018), Winogrande (Sakaguchi et al., 2021), and PIQA (Bisk et al., 2020). These benchmarks are structured in a multiple-choice format. We used each question along with its correct answer as input to compute the respective evaluation metrics.
Prompt Learning Datasets. We conducted prompt learning experiments using the OpenOrca dataset (Lian et al., 2023b), where each question-answer (QA) pair features a unique system prompt. To explore the impact of these prompts on Matrix Nuclear-Norm, we carefully selected three generally instructive prompts (see Table 13) to observe their roles in the QA pairs.
Validation Datasets. Additionally, we performed inference and evaluation on two widely used datasets: AlpacaEval (Dubois et al., 2024) and Chatbot Arena (Chiang et al., 2024). We conducted inference on both datasets, evaluated the responses using our proposed metrics, and ranked the models accordingly.
5.2 Matrix Nuclear-Norm Observation
5.2.1 A Comparative Analysis of Computational Time
Figure 1: CEREBRAS-GPT: Time comparison
To evaluate the computational efficiency of Matrix Nuclear-Norm in comparison to Matrix Entropy for LLMs, we conducted experiments across various model sizes using multiple benchmark datasets. The results, summarized in Table 1, demonstrate a clear advantage of Matrix Nuclear-Norm in terms of computation time, particularly for larger models.
As model sizes increased, Matrix Entropy’s computation time rose dramatically, reaching approximately 16.3 hours for the 13B model . In contrast, Matrix Nuclear-Norm only required about 0.82 hours for the same model, representing nearly a 20-fold reduction in computation time. This trend was consistent across all model sizes, with Matrix Nuclear-Norm consistently proving to be much faster (as illustrated in Figure 1). For example, the  [^1] M model showed that Matrix Nuclear-Norm was 8.58 times quicker than Matrix Entropy.
The significant efficiency gain is attributed to the lower computational complexity of Matrix Nuclear-Norm, $O(m\cdot n+n\log n)$, compared to Matrix Entropy’s $O(n^{3})$. This makes Matrix Nuclear-Norm not only an effective but also a highly efficient metric for evaluating LLMs, especially as models continue to scale in size.
In summary, Matrix Nuclear-Norm achieves comparable evaluation accuracy to Matrix Entropy but with vastly superior computational efficiency, making it a practical and scalable choice for assessing LLMs.
Model Size Matrix Entropy Time (s) Matrix Nuclear-Norm Time (s) Ratio
 [^1] M 623.5367 72.6734 8.5800
256M 1213.0604 110.8692 10.9414
590M 2959.6949 184.7785 16.0175
1.3B 6760.1893 379.0093 17.8365
2.7B 12083.7105 732.6385 16.4934
6.7B 38791.2035 1598.4151 24.2685
13B 59028.4483 2984.1529 19.7806
Table 1: CEREBRAS-GPT: Time Comparison between Matrix Entropy and Matrix Nuclear-Norm
5.2.2 Scaling Law of Matrix Nuclear-Norm
To affirm Matrix Nuclear-Norm’s efficacy as an evaluative metric, we evaluated Cerebras-GPT models on four datasets including dolly-15k, Wikipedia, openwebtext2, and hh-rlhf comparing Matrix Nuclear-Norm, matrix entropy, perplexity, and loss. Results, detailed in Table 11 (Appendix), demonstrate Matrix Nuclear-Norm’s consistent decrease with model size enlargement, signifying better data compression and information processing in larger models. This trend (see in Figure 2(d)) validates Matrix Nuclear-Norm’s utility across the evaluated datasets. Notably, anomalies at the 2.7B and 13B highlight areas needing further exploration.
(a) Cross-Entropy Loss
(b) Perplexity
(c) Matrix Entropy
(d) Matrix Nuclear-Norm
Figure 2: Comparison of Matrix Nuclear-Norm, matrix entropy, loss, and perplexity when model scales up.
5.2.3 Relationship of Benchmark INDICATORS
Findings indicate the efficacy of the Matrix Nuclear-Norm as a metric for evaluating LLM, as shown in Table 10 (Appendix), there is an overall downward trend in Matrix Nuclear-Norm values with increasing model sizes, signifying enhanced compression efficiency. However, notable anomalies at the 2.7B and 13B checkpoints suggest that these specific model sizes warrant closer examination. Despite these discrepancies, the Matrix Nuclear-Norm consistently demonstrates superior computational efficiency and accuracy compared to traditional metrics, highlighting its promising applicability for future model evaluations.
5.3 Language Investigation
5.3.1 Sentence Operation Experiments
Figure 3 clearly indicates that sentence manipulations significantly influence Matrix Nuclear-Norm values, which generally decline as model size increases. This trend confirms the enhanced information compression capabilities of larger models. The ranking of Matrix Nuclear-Norm values by operation is as follows: Reverse $>$ Shuffle & Reverse $>$ Shuffle $>$ Base. This indicates that disrupting sentence structure through Reverse and Shuffle & Reverse operations leads to higher Matrix Nuclear-Norm values due to increased information chaos and processing complexity. In contrast, the Shuffle operation has minimal effect on compression, while the Base condition consistently yields the lowest Matrix Nuclear-Norm values, signifying optimal information compression efficiency with unaltered sentences.
Despite the overall downward trend in Matrix Nuclear-Norm values with increasing model size, the 2.7B model exhibits slightly higher values for Shuffle and Base operations compared to the 1.3B model. This anomaly suggests that the 2.7B model may retain more nuanced information when handling shuffled data or operate through more intricate mechanisms. However, this does not detract from the overarching conclusion that larger models excel at compressing information, thereby demonstrating superior processing capabilities.
Figure 3: Results of sentence operation. Shuffling and reversing disrupt the text structure and diminish the informational content, leading to an increase in Matrix Nuclear-Norm.
5.3.2 Analysis of Length Dynamics
Figure 4: The Matrix Nuclear-Norm values for contexts of varying lengths show that as text length increases, the Matrix Nuclear-Norm continues to rise and tends to converge.
The analysis reveals that Matrix Nuclear-Norm values generally increase as input length rises, aligning with our expectations (see Figure 4). Longer inputs necessitate that the model manage and compress more information, which naturally leads to higher Matrix Nuclear-Norm values. Most models exhibit this trend, indicating effective handling of the increased information load.
However, the gpt-2.7B and gpt-13B models display anomalies in their Matrix Nuclear-Norm values at 64 and 128 tokens, where the value at 128 tokens is lower than that at 64 tokens. This discrepancy may be attributed to these models employing different information compression mechanisms or optimization strategies tailored to specific input lengths, allowing for more effective compression at those lengths.
Overall, aside from a few outliers, the results largely conform to expectations, demonstrating that Matrix Nuclear-Norm values increase with input length, reflecting the greater volume and complexity of information that models must handle.
To address the observed trend of rising Matrix Nuclear-Norm values with longer sentences, we incorporated a normalization step in our methodology via dividing the Matrix Nuclear-Norm values by the sentence length. This adjustment helps mitigate any biases introduced by models that tend to generate longer sentences during inference.
GPT MODEL SIZE
LENGTH  [^1] M 256M 590M 1.3B 2.7B 6.7B 13B
64 0.4574 0.4125 0.3787 0.3486 0.4053 0.3315 0.4148
128 0.5293 0.4680 0.4270 0.3835 0.4143 0.3477 0.4032
512 0.7883 0.6978 0.6251 0.5554 0.5265 0.4468 0.4422
1024 0.9132 0.8787 0.7802 0.6953 0.6351 0.5383 0.5028
Table 2: Analysis of Length Dynamics
5.3.3 Analysis of Prompt Learning
The experimental results (shown in Table 3) indicate that we performed inference on different sizes of GPT models using three carefully selected prompts (shown in Table 13) and calculated the Matrix Nuclear-Norm values of their responses. As the model size increased, the Matrix Nuclear-Norm values gradually decreased, demonstrating that larger models possess greater information compression capabilities. The prompts significantly influenced Matrix Nuclear-Norm, with variations reflecting the models’ responses to prompt complexity. Specifically, GPT-1.3B showed a notable decrease in Matrix Nuclear-Norm after the input prompts, indicating its sensitivity to them, while GPT-2.7B exhibited smaller changes. In contrast, GPT-6.7B displayed minimal variation across all prompts, suggesting stable performance regardless of prompt detail. Overall, more detailed prompts resulted in larger information volumes in the model’s responses, leading to corresponding changes in Matrix Nuclear-Norm values.
Table 3: Results of prompt learning with (Empty Prompt) and without (Prompt 1, 2, 3) the use of prompts. Incorporating prompts as prefixes before the QA pairs enhances the models’ ability to achieve better compression.
ADDING PROMPT TO QA PAIRS
MODELS EMPTY PROMPT PROMPT 1 PROMPT 2 PROMPT 3 AVERAGE $\Delta x$
CEREBRAS-GPT-1.3B 0.150955 0.147577 0.140511 0.141358 0.14453 ↓0.006425
CEREBRAS-GPT-2.7B 0.150130 0.151522 0.142834 0.151842 0.14844 ↓0.001690
CEREBRAS-GPT-6.7B 0.132042 0.128346 0.124094 0.133211 0.12923 ↓0.002812
6 Implementing Proposed Metrics: Evaluating and Ranking Language Models in Practice
6.1 Inference-Based Model Assessment
In this section, we evaluated model inference across the AlpacaEval and Chatbot Arena benchmarks using the Matrix Nuclear-Norm metric prior to the final MLP classification head. The analysis revealed that Matrix Nuclear-Norm reliably ranks model performance, with lower values indicating enhanced information processing efficiency, particularly as model size scales up.
For instance, the Llama-3 70B model demonstrated superior compression capabilities compared to its 8B counterpart, as reflected by significantly lower Matrix Nuclear-Norm values across both benchmarks (see Table A.2 in the Appendix). A similar trend was observed in the Vicuna family, where Matrix Nuclear-Norm values consistently decreased from 0.4623 for the 7B model to 0.3643 for the 33B model on the AlpacaEval dataset, indicating progressive improvements in information handling (see Table 4). Additionally, the DeepSeek models exhibited a consistent decrease in Matrix Nuclear-Norm values as model size increased, further demonstrating the metric’s validity.
Overall, these results substantiate Matrix Nuclear-Norm as a robust and reliable tool for evaluating and ranking LLMs, demonstrating its capacity to capture critical aspects of model performance across diverse benchmarks.
Model DataSet 7B 13B 33B
Vicuna Alpaca 0.4623 0.4159 0.3643
Arena 0.4824 0.4311 0.3734
Model DataSet 1.3B 6.7B 7B
DeepSeek Alpaca 0.4882 0.3472 0.3352
Arena 0.5754 0.4175 0.4357
Table 4: Matrix Nuclear-Norms in Vicuna and DeepSeek Responses
6.2 Matrix Nuclear-Norm Benchmarking: Ranking Mid-Sized Models
In this experimental section, we utilized Matrix Nuclear-Norm to evaluate the responses of LLMs, focusing on 7B and 70B variants. Notably, lower Matrix Nuclear-Norm values indicate more efficient information compression, serving as a robust indicator of model performance.
Among the 7B models, DeepSeek-7B exhibited the most efficient information processing with the lowest average Matrix Nuclear-Norm score of 0.3855 across Alpaca and Arena datasets (see Table 4). Gemma-7B followed closely with an average score of 0.3879, whereas QWEN 2-7B demonstrated less efficient compression with an average score of 0.5870. In contrast, the 70B models showed varied performance, with Llama 2-70B achieving the best average score of 0.3974, slightly outperforming Llama 3-70B (0.4951) and QWEN models, which scored around 0.5.
Interestingly, certain 7B models, like DeepSeek-7B and Gemma-7B, outperformed larger 70B models, underscoring that model efficiency is not solely determined by size. These results highlight that factors such as architecture, training methodology, and data complexity play crucial roles in information processing capabilities beyond scale.
MODEL Matrix Nuclear-Norm Rank
Alpaca Arena-Hard Avg Score
DeepSeek-7B 0.3352 0.4357 0.3855 $\downarrow$
Gemma-7B 0.3759 0.3998 0.3879 $\downarrow$
Vicuna-7B 0.4623 0.4824 0.4724 $\downarrow$
LLaMA 2-7B 0.4648 0.5038 0.4843 $\downarrow$
QWEN 1.5-7B 0.4866 0.5165 0.5016 $\downarrow$
Mistral-7B 0.4980 0.5126 0.5053 $\downarrow$
QWEN 2-7B 0.5989 0.5751 0.5870 $\downarrow$
QWEN 1.5-72B 0.5291 0.5065 0.5178 $\downarrow$
QWEN 2-72B 0.5261 0.4689 0.4975 $\downarrow$
Llama 3-70B 0.4935 0.4967 0.4951 $\downarrow$
Llama 2-70B 0.3862 0.4086 0.3974 $\downarrow$
Table 5: Matrix Nuclear-Norm Rankings: A Comparative Analysis of Model Performance
To validate the design rationale and robustness of the Matrix Nuclear-Norm, we conducted a series of ablation studies. Due to space constraints, detailed results are provided in A.1 (appendix) to maintain brevity in the main text. These experiments included evaluations across different model families, such as Cerebras-GPT and Pythia, as well as comparisons of various data sampling strategies.The results demonstrate that the Matrix Nuclear-Norm consistently performs well across different model scales and sampling variations. This not only confirms its applicability across diverse models but also verifies its stability and reliability in handling large-scale datasets. We also provide an ablation study in the appendix, further proving the method’s efficiency and accuracy in evaluating LLMs.
7 Conclusion
In conclusion, Matrix Nuclear-Norm stands out as a promising evaluation metric for LLMs, offering significant advantages in assessing information compression and redundancy elimination. Its key strengths include remarkable computational efficiency, greatly exceeding that of existing metrics like matrix entropy, along with exceptional stability across diverse datasets. Matrix Nuclear-Norm’s responsiveness to model performance under varying inputs emphasizes its ability to gauge not only performance but also the intricate adaptability of models. This metric marks a significant advancement in NLP, establishing a clear and effective framework for future research and development in the evaluation and optimization of language models.
8 Limitations
Although Matrix Nuclear-Norm performs well in evaluating the performance of LLMs, it still has some limitations. First, since Matrix Nuclear-Norm’s computation relies on the model’s hidden states, the evaluation results are sensitive to both the model architecture and the training process. As a result, under different model designs or training settings, especially for models like GPT-1.3B and GPT-2.7B, inconsistencies in Matrix Nuclear-Norm’s performance may arise, limiting its applicability across a wider range of models. Additionally, while Matrix Nuclear-Norm offers computational efficiency advantages over traditional methods, it may still face challenges with resource consumption when evaluating extremely large models. As model sizes continue to grow, further optimization of Matrix Nuclear-Norm’s computational efficiency and evaluation stability is required.
9 Ethics Statement
Our study adheres to strict ethical guidelines by utilizing only publicly available and open-source datasets. We ensured that all datasets used, such as dolly-15k, hh-rlhf, OpenBookQA, Winogrande, PIQA, AlpacaEval, and Chatbot Arena, are free from harmful, biased, or sensitive content. Additionally, careful curation was conducted to avoid toxic, inappropriate, or ethically problematic data, thereby ensuring the integrity and safety of our research. This commitment reflects our dedication to responsible AI research and the broader implications of using such data in language model development.
Appendix A Appendix
A.1 Ablation Study
To thoroughly validate the rationale behind our metric design, experimental framework, and the efficacy of Matrix Nuclear-Norm, we conducted a series of ablation studies.
A.1.1 Different Model Family
In addition to evaluating Matrix Nuclear-Norm within the Cerebras-GPT model series, we extended our experiments to the Pythia model family, which spans from 14M to 12B parameters and is trained on consistent public datasets. Utilizing the same datasets as described in Section 5.2.2, we computed matrix entropy, loss values, and Matrix Nuclear-Norm for these models. The empirical results (see Figure 5(c)) demonstrate that the Matrix Nuclear-Norm values for the Pythia models adhere to established scaling laws. However, we excluded metrics for the 14M, 31M, and 1B models due to notable deviations from the expected range, likely stemming from the inherent instability associated with smaller parameter sizes when tackling complex tasks. This further reinforces Matrix Nuclear-Norm as a robust metric for assessing model performance, underscoring its utility in the comparative analysis of LLMs.
Moreover, we compared the computation times for Matrix Entropy and Matrix Nuclear-Norm across the Pythia models (can see in Figure 7). The results unequivocally indicate that Matrix Nuclear-Norm necessitates considerably less computation time than Matrix Entropy, underscoring its efficiency. Detailed results are summarized in Table 12.
(a) Cross-Entropy Loss
(b) Matrix Entropy
(c) Matrix Nuclear-Norm
Figure 5: Pythia Model Metrics: Matrix Nuclear-Norm, Matrix Entropy, and Loss
A.1.2 Sampling Strategy
In the ablation experiments, we extracted a baseline subset of 10,000 entries from the extensive Wikipedia dataset using three random seeds to evaluate the robustness of the Matrix Nuclear-Norm metric. We also tested additional subsets of 15,000 and 20,000 entries due to potential entry count issues. Given the large scale of the datasets, comprehensive calculations were impractical, so we employed random sampling.
The results showed that variations in random seeds and sample sizes had minimal impact on Matrix Nuclear-Norm values, with a standard deviation of only 0.0004975 (see Table 6), indicating high consistency across trials. These findings confirm the Matrix Nuclear-Norm as a reliable metric for large-scale datasets, effectively evaluating information compression and redundancy elimination in LLMs.
Table 6: Ablation study of differnet sampling strategies on the Wikimedia[18] dataset.
MODEL SAMPLING STRATEGY STANDARD DEVIATION
10000 (SEED 1) 10000 (SEED 2) 10000 (SEED 3) 15000 20000
CEREBRAS-GPT-1.3B 0.5684 0.5670 0.5676 0.5699 0.5693 0.0004975
A.2 Supplementary Experiment Results
The following results provide additional insights into the Matrix Nuclear-Norm evaluations and comparisons across various language models:
  1. 1.
    Tables A.2 and A.2 present the Matrix Nuclear-Norm evaluation results during the inference process for Llama-3 and QWEN-2.
  2. 2.
    Figure 6 illustrates that as model size increases, the computation time for Matrix Entropy grows exponentially, while Matrix Nuclear-Norm demonstrates a significant time advantage. This further emphasizes Matrix Nuclear-Norm’s efficiency in assessing model performance.The complete results are presented in Table 7, which includes all relevant time data for the Pythia model family.
  3. 3.
    Table 11 contains the complete results for the comparison of Matrix Nuclear-Norm and other metrics based on Cerebras-GPT family considered in Figure 2(d).
  4. 4.
    Table 10 demonstrates the correlation between Matrix Nuclear-Norm and other benchmark indicators, showing a consistent trend where values decrease as model size increases. This analysis examines the performance of language modeling indicators across OpenBookQA, Winogrande, and PIQA datasets.
  5. 5.
    Table 12 illustrates the numerical results of Figure 5(c) in the ablation study of Pythia family.
  6. 6.
    Table 13 shows the prompts used for the investigation of prompt learning.
Figure 6: Pythia: Time Comparison of Matrix Entropy and Nuclear-Norm
Model Size Matrix Entropy Time (s) Matrix Nuclear-Norm Time (s) Ratio
14M 52.8669 22.2652 2.3772
31M 114.0820 28.1842 4.0477
70M 320.6641 24.3188 13.1855
160M 631.9762 41.6187 15.1817
410M 1040.9764 80.9814 12.8481
1B 4650.1264 114.0639 40.8387
1.4B 6387.0392 347.8670 18.3858
2.8B 8127.1343 343.3888 23.6778
6.9B 28197.8172 816.6332 34.5350
12B 47273.5235 1276.1128 37.0485
Table 7: Pythia Model: Matrix Entropy vs. Matrix Nuclear-Norm Time Comparison
Model DataSet 0.5B 1.5B 7B 72B
Alpaca 0.6551 0.6176 0.5989 0.5261
QWEN 2 Arena 0.6872 0.6374 0.5751 0.4689
Table 8: Matrix Nuclear-Norm in QWEN 2 Responses
Model Data Set 8B 70B
Llama-3 Alpaca 0.5782 0.4935
Arena 0.5817 0.4967
Table 9: Matrix Nuclear-Norm in Llama3 esponses
GPT MODEL SIZE
BENCHMARKS INDICATORS  [^1] M 256M 590M 1.3B 2.7B 6.7B 13B
ACCURACY 0.118 0.158 0.158 0.166 0.206 0.238 0.286
MATRIX ENTROPY 0.3575 0.3416 0.3237 0.3140 0.2991 0.2848 0.2767
LOSS 5.6196 5.3536 5.1881 4.9690 4.8723 4.7195 4.7050
PPL 148.38 108.10 83.45 65.10 50.93 41.80 40.89
OPENBOOKQA MATRIX NUCLEAR-NORM 0.4447 0.4057 0.3941 0.3644 0.4606 0.3672 0.4423
ACCURACY 0.488 0.511 0.498 0.521 0.559 0.602 0.646
MATRIX ENTROPY 0.4073 0.3915 0.3706 0.3605 0.3419 0.3272 0.3149
LOSS 4.7869 4.5854 4.4141 4.2513 4.1107 4.0109 4.0266
PPL 39.81 30.25 26.57 21.87 18.55 16.53 16.94
WINOGRANDE MATRIX NUCLEAR-NORM 0.4802 0.4479 0.4440 0.4133 0.5232 0.4220 0.4964
ACCURACY 0.594 0.613 0.627 0.664 0.701 0.739 0.766
MATRIX ENTROPY 0.4168 0.3991 0.3783 0.3676 0.3504 0.3344 0.3264
LOSS 4.8425 4.5470 4.4029 4.1613 4.0075 3.8545 3.8826
PPL 69.80 47.94 37.88 28.76 23.15 19.76 19.72
PIQA MATRIX NUCLEAR-NORM 0.4868 0.4327 0.4164 0.3826 0.4452 0.3675 0.4149
Table 10: Language modeling indicators on openbookqa, winogrande and piqa.
DATASET INDICATORS GPT MODELS SIZE
 [^1] M 256M 590M 1.3B 2.7B 6.7B 13B
DOLLY-15K MATRIX ENTROPY 0.5976 0.5840 0.5582 0.5477 0.5240 0.5064 0.4859
LOSS 3.6710 3.2907 3.0359 2.7517 2.5015 2.2911 2.3098
PPL 39.93 27.53 21.42 16.15 12.50 10.23 10.30
MATRIX NUCLEAR-NORM 0.6207 0.5565 0.5063 0.4553 0.4639 0.3904 0.4859
WIKIPEDIA MATRIX ENTROPY 0.6177 0.6077 0.5848 0.5786 0.5523 0.5368 0.5126
LOSS 3.2900 2.9343 2.6854 2.4282 2.2045 2.0216 2.0327
PPL 31.38 22.51 17.89 13.85 11.08 9.19 9.32
MATRIX NUCLEAR-NORM 0.6744 0.6422 0.6094 0.5639 0.5438 0.4660 0.4708
OPENWEBTEXT2 MATRIX ENTROPY 0.6527 0.6479 0.6206 0.6142 0.5855 0.5683 0.5463
LOSS 3.7509 3.3852 3.1414 2.8860 2.6465 2.4708 2.4685
PPL 36.79 25.82 20.34 15.89 12.51 10.57 10.51
MATRIX NUCLEAR-NORM 0.7147 0.7066 0.6823 0.6363 0.6017 0.5133 0.4991
HH-RLHF MATRIX ENTROPY 0.5753 0.5635 0.5350 0.5268 0.4971 0.4813 0.4640
LOSS 3.3078 2.9964 2.8171 2.6431 2.4622 2.3526 2.3323
PPL 18.97 14.01 11.62 9.73 8.12 7.27 7.19
MATRIX NUCLEAR-NORM 0.6309 0.5716 0.5307 0.4771 0.4959 0.4277 0.4518
Table 11: Language modeling indicators on the Dolly-15k, Wikipedia, OpenWebText2, and HH-RLHF datasets.
Table 12: The language modeling indicators (where lower values indicate better performance) for the Pythia models were evaluated on the Dolly-15k, Wikipedia, OpenWebText2, and HH-RLHF datasets.
PYTHIA MODELS SIZE
DATASETS INDICATORS 14M 31M 70M 160M 410M 1B 1.4B 2.8B 6.9B 12B
MATRIX ENTROPY 0.7732 0.7155 0.6707 0.6243 0.5760 0.5328 0.5309 0.5263 0.5003 0.4876
LOSS 4.4546 4.0358 3.5990 3.1323 2.6752 2.4843 2.3816 2.2484 2.1368 2.0616
DOLLY-15K MATRIX NUCLEAR-NORM 0.7508 0.7735 0.6984 0.6104 0.5760 0.4710 0.4922 0.4585 0.4202 0.4181
MATRIX ENTROPY 0.7938 0.7442 0.7003 0.6580 0.6039 0.5584 0.5587 0.5553 0.5314 0.5140
LOSS 4.1112 3.6921 3.2694 2.8207 2.4017 2.2213 2.1292 2.0140 1.9120 1.8489
WIKIPEDIA MATRIX NUCLEAR-NORM 0.6053 0.6700 0.6996 0.6718 0.6464 0.5591 0.5787 0.5410 0.4850 0.4768
MATRIX ENTROPY 0.8144 0.7749 0.7370 0.6980 0.6415 0.5944 0.5916 0.5887 0.5591 0.5417
LOSS 4.3965 4.0033 3.6284 3.2031 2.7838 2.6198 2.5228 2.4005 2.3133 2.2502
OPENWEBTEXT2 MATRIX NUCLEAR-NORM 0.5041 0.6186 0.7142 0.7258 0.7105 0.6215 0.6378 0.5967 0.5275 0.5110
MATRIX ENTROPY 0.7673 0.7114 0.6607 0.6126 0.5552 0.5054 0.5032 0.4977 0.4699 0.4528
LOSS 3.7466 3.4018 3.1146 2.7366 2.4340 2.3311 2.2687 2.1992 2.1199 2.0905
HH-RLHF MATRIX NUCLEAR-NORM 0.7353 0.7674 0.6964 0.6182 0.5886 0.4825 0.5141 0.4839 0.4562 0.4481
Table 13: The prompts selected from OpenOrca[30] dataset.
Prompt ID Prompt Content
Prompt 1 You are an AI assistant. You will be given a task. You must generate a detailed and long answer.
Prompt 2 You are a helpful assistant, who always provide explanation. Think like you are answering to a five year old.
Prompt 3 You are an AI assistant. User will give you a task. Your goal is to complete the task as faithfully as you can. While performing the task think step-by-step and justify your steps.
A.3 Analysis of Algorithmic Complexity
The primary computational expense of Matrix Nuclear-Norm arises from the calculation and sorting of the L2 norm of the matrix. By avoiding Singular Value Decomposition (SVD), we reduce the time complexity from the traditional nuclear norm of $O(n^{3})$ to $O(n^{2})$, giving Matrix Nuclear-Norm a significant advantage in handling large-scale data. This reduction in complexity greatly enhances the algorithm’s practicality, especially for applications involving large matrices.
When analyzing the time complexity of the newly proposed Matrix Nuclear-Norm (L2-Norm Based Approximation of Nuclear Norm) against traditional Matrix Entropy, our objective is to demonstrate that Matrix Nuclear-Norm significantly outperforms Matrix Entropy in terms of time efficiency. We will support this claim with detailed complexity analysis and experimental results.
A.3.1 Time Complexity Analysis
Analysis 1: Time Complexity of Matrix Entropy
The computation of Matrix Entropy involves several complex steps, with the key bottleneck being Singular Value Decomposition (SVD), which is central to computing eigenvalues. The following steps primarily contribute to the time complexity:
  1. 1.
    Matrix Normalization: This step has a time complexity of $O(m\cdot n)$, where $m$ is the number of rows and $n$ is the number of columns.
  2. 2.
    Computing the Inner Product Matrix: Calculating $Z^{T}Z$ has a time complexity of $O(n^{2}\cdot m)$ due to the multiplication of two matrices sized $m\times n$.
  3. 3.
    Singular Value Decomposition (SVD): The time complexity of SVD is $O(n^{3})$, which is the primary computational bottleneck, especially for large $n$.
Therefore, the total time complexity of Matrix Entropy can be approximated as:
$O(m\cdot n+n^{2}\cdot m+n^{3})=O(n^{3})$
This complexity indicates that Matrix Entropy becomes increasingly impractical for large-scale models as $n$ grows.
Analysis 2: Time Complexity of Matrix Nuclear-Norm
Matrix Nuclear-Norm avoids the SVD step by approximating the nuclear norm using the L2 norm, resulting in a more efficient computation. The analysis is as follows:
  1. 1.
    Matrix Normalization: Similar to Matrix Entropy, this step has a time complexity of $O(m\cdot n)$.
  2. 2.
    Calculating the L2 Norm: For each column vector, the L2 norm is computed with a complexity of $O(m\cdot n)$, where we take the square root of the sum of squares for each column vector.
  3. 3.
    Sorting and Extracting the Top D Features: Sorting the L2 norms has a complexity of $O(n\log n)$.
Therefore, the overall time complexity of Matrix Nuclear-Norm is:
$O(m\cdot n+n\log n)\approx O(n^{2})\quad\text{when}\quad m\approx n$
This indicates that Matrix Nuclear-Norm is computationally more efficient, especially as $n$ increases.
A.3.2 Experimental Validation and Comparative Analysis
To empirically validate the theoretical time complexities, we conducted experiments using matrices of various sizes. Figure 6 shows that as $n$ increases, Matrix Nuclear-Norm consistently outperforms Matrix Entropy in terms of runtime, confirming the theoretical advantage.
Discussion of Assumptions and Applicability Our complexity analysis assumes $m\approx n$, which holds in many real-world applications, such as evaluating square matrices in large-scale language models. However, in cases where $m\neq n$, the time complexity might differ slightly. Nonetheless, Matrix Nuclear-Norm is expected to maintain its efficiency advantage due to its avoidance of the costly SVD operation.
Impact of Constant Factors Although both $O(n^{2})$ and $O(n^{3})$ indicate asymptotic behavior, Matrix Nuclear-Norm’s significantly smaller constant factors make it computationally favorable even for moderately sized matrices, as evidenced in our experimental results.
A.3.3 Conclusion of the Complexity Analysis
Through this detailed analysis and experimental validation, we conclude the following:
  • Matrix Entropy, with its reliance on SVD, has a time complexity of $O(n^{3})$, making it computationally expensive for large-scale applications.
  • Matrix Nuclear-Norm, by using the L2 norm approximation, achieves a time complexity of $O(m\cdot n+n\log n)\approx O(n^{2})$, significantly reducing computational costs.
  • Experimental results confirm that Matrix Nuclear-Norm offers superior time efficiency for evaluating large-scale models, particularly those with millions or billions of parameters.
A.4 Proof of Theorem 1: Detailed Version
Here we present the proof of Theorem 1, demonstrating the strictly inverse monotonic relationship between $H(A)$ and $\|A\|_{F}$.
Given a matrix $A\in\mathbb{R}^{B\times C}$, where each element is non-negative and each row sums to 1, i.e.,
$\sum_{j=1}^{C}A_{i,j}=1\quad\text{for}\ i=1,2,\ldots,B.$
We define the Shannon entropy $H(A)$ as:
$H(A)=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{C}A_{i,j}\log(A_{i,j}),$
and the Frobenius norm $\|A\|_{F}$ as:
$\|A\|_{F}=\sqrt{\sum_{i=1}^{B}\sum_{j=1}^{C}|A_{i,j}|^{2}}.$
Step 1: Analysis for a Fixed Row Vector
For a fixed row $i$, let $a_{j}=A_{i,j}$ such that $a_{j}\geq 0$ and $\sum_{j=1}^{C}a_{j}=1$. The entropy for this single row is given by:
$H_{i}=-\sum_{j=1}^{C}a_{j}\log a_{j},$
and the Frobenius norm for this row vector $a$ is:
$\|a\|_{2}=\sqrt{\sum_{j=1}^{C}a_{j}^{2}}.$
To determine when $H_{i}$ is maximized or minimized, we apply the method of Lagrange multipliers. Consider the function:
$L(a_{1},a_{2},\ldots,a_{C},\lambda)=-\sum_{j=1}^{C}a_{j}\log a_{j}+\lambda\left(\sum_{j=1}^{C}a_{j}-1\right).$
Taking partial derivatives with respect to $a_{j}$ and setting them to zero:
$\frac{\partial L}{\partial a_{j}}=-\log a_{j}-1+\lambda=0\implies a_{j}=e^{\lambda-1}.$
Since $\sum_{j=1}^{C}a_{j}=1$, we find $a_{j}=\frac{1}{C}$ for all $j$. Thus, $H_{i}$ is maximized when all $a_{j}$ are equal. In this case,
$\|a\|_{2}=\sqrt{\sum_{j=1}^{C}\left(\frac{1}{C}\right)^{2}}=\sqrt{\frac{1}{C}}.$
Conversely, $H_{i}$ reaches its minimum when one element $a_{k}=1$ and all other $a_{j}=0$. At this point,
$\|a\|_{2}=\sqrt{1^{2}+0}=1.$
This demonstrates that $H_{i}$ and $\|a\|_{2}$ are inversely related for each row.
Step 2: Inverse Monotonicity for the Entire Matrix
Since $H(A)$ is the average of all $H_{i}$ values:
$H(A)=\frac{1}{B}\sum_{i=1}^{B}H_{i},$
and $\|A\|_{F}$ is the square root of the sum of the squared norms $\|a\|_{2}^{2}$ for each row:
$\|A\|_{F}=\sqrt{\sum_{i=1}^{B}\|a\|_{2}^{2}},$
the inverse relationship between $H_{i}$ and $\|a\|_{2}$ implies that $H(A)$ and $\|A\|_{F}$ are also inversely related across the entire matrix $A$.
Step 3: Determining the Bounds of $\|A\|_{F}$
To find the bounds, first note that when $a_{j}=\frac{1}{C}$ for all $j$,
$\|A\|_{F}=\sqrt{\sum_{i=1}^{B}\sum_{j=1}^{C}\left(\frac{1}{C}\right)^{2}}=\sqrt{B\cdot\frac{1}{C}}=\sqrt{\frac{B}{C}},$
which corresponds to the minimum discriminability.
For the upper bound, when each row has only one element equal to 1,
$\|A\|_{F}=\sqrt{\sum_{i=1}^{B}1}=\sqrt{B},$
indicating maximum discriminability.
Step 4: Practical Implications
The inverse monotonic relationship between $H(A)$ and $\|A\|_{F}$ is significant in machine learning and classification tasks, where maximizing $\|A\|_{F}$ implies achieving greater certainty in predictions. This property can be used to evaluate model performance, where higher $\|A\|_{F}$ corresponds to a model that is more confident and discriminative in its predictions. Therefore, $H(A)$ and $\|A\|_{F}$ are strictly inversely monotonic, and the proof is complete.
A.5 Proof of Theorem 3
Given that $\|A\|_{F}$ is near $\sqrt{B}$, we want to approximate the $j$-th largest singular value $\sigma_{j}$ as $\text{top}\left(\sum_{i=1}^{B}A_{i,j}^{2},j\right)$. This approximation will be derived by examining the contributions of the columns of $A$.
1. Decomposition of $A$ and Eigenvalues: By singular value decomposition (SVD), $A=U\Sigma V^{T}$, where $\Sigma$ is a diagonal matrix containing singular values $\sigma_{1},\sigma_{2},\ldots,\sigma_{D}$.
When looking at the product $A^{T}A$, the element at position $(j,j)$ is:
$(A^{T}A)_{j,j}=\sum_{i=1}^{B}A_{i,j}^{2}$
This sum represents the total squared magnitude of the entries in the $j$-th column of $A$.
2. Connecting to Singular Values: Since singular values represent the contribution of orthogonal projections of $A$, if $\|A\|_{F}$ is near $\sqrt{B}$, the main contribution to the nuclear norm $\|A\|_{*}$ comes from columns with the largest sums of squared values $\sum_{i=1}^{B}A_{i,j}^{2}$.
3. Approximation: For matrices with well-distributed entries in each column (a common property when $\|A\|_{F}\approx\sqrt{B}$), the top singular values can be effectively estimated by sorting these column norms. Hence, the largest singular values $\sigma_{j}$ approximately correspond to the largest column norms $\sqrt{\sum_{i=1}^{B}A_{i,j}^{2}}$.
Therefore, for each $j\in\{1,\ldots,D\}$,
$\sigma_{j}\approx\text{top}\left(\sum_{i=1}^{B}A_{i,j}^{2},j\right).$
This concludes the proof that the $j$-th largest singular value $\sigma_{j}$ can be approximated by the top $j$ values of column-wise norms of $A$, as stated in Theorem 3.

### v3

@@
 […]
+Large Language Model Evaluation via Matrix Nuclear-Norm
+Yahan Li1, Tingyu Xia1, Yi Chang1,2,3, Yuan Wu1,2thanks: Corresponding author 1School of Artificial Intelligence, Jilin University 2Key Laboratory of Symbolic Computation and Knowledge Engineering, Jilin University 3International Center of Future Science, Jilin University yahan23@mails.jlu.edu.cn, xiaty21@mails.jlu.edu.cn, yichang@jlu.edu.cn yuanwu@jlu.edu.cn
 1 Introduction
-Large language models (LLMs), such as Gemini (Gemini et al., 2023), Llama (Touvron et al., 2023), and GPT-4 (GPT-4 Achiam et al., 2023), have demonstrated remarkable performance across a variety of natural language processing (NLP) tasks (Zhao et al., 2023).
-They are not only transforming the landscape of NLP (Saul et al., 2005; Liu et al., 2023; Sawada et al., 2023) but also bring beneficial impacts on computer vision (Lian et al., 2023a; Wang et al., 2024) and graph neural networks (Zhang et al., 2024; Chen et al., 2024b), achieving stellar results on various leaderboards.
-Despite these advancements, assessing a model’s ability to compress information remains a crucial research challenge (Delétang et al., 2023).
-Compression focuses on efficiently distilling essential information from vast training datasets while discarding redundant elements, showcasing a model’s ability to learn and recognize the underlying structure of the data (Wei et al., 2024).
-LLMs are expected to perform this form of compression during their training process (Zhao et al., 2023).
-Specifically, in the early stages of training, after random initialization, the representations produced from the data are often chaotic.
-However, as training progresses, these representations become more organized, allowing the model to filter out unnecessary information.
-Hence, assessing an LLM’s capacity for information compression is crucial for understanding its learning efficiency and representational power.
-Current compression evaluation methods, such as Matrix Entropy introduced by Wei et al.
-(2024), measure information compression efficiency through processing models’ output representations on datasets.
-However, the reliance of Matrix Entropy on Singular Value Decomposition (SVD) (Kung et al., 1983; Zhang, 2015) leads to significant computational complexity, typically $O(n^{3})$, which limits its applicability in large-scale models.
+Large language models (LLMs), such as Gemini (Gemini et al., 2023), Deepseek (Guo et al., 2025), and GPT-4 (GPT-4 Achiam et al., 2023), have shown exceptional performance in numerous natural language processing (NLP) tasks (Zhao et al., 2023).
+These models are transforming the way we approach NLP tasks, providing unprecedented capabilities and solutions to complex problems.
+They are revolutionizing NLP (Saul et al., 2005; Liu et al., 2023; Sawada et al., 2023) and positively impacting computer vision (Lian et al., 2023a; Wang et al., 2024) and graph neural networks (Zhang et al., 2024; Chen et al., 2024), achieving top results on leaderboards.
+Despite these advancements, evaluating a model’s ability to compress information remains a critical research challenge (Delétang et al., 2023).
+This challenge is essential for improving the overall efficiency of these models.
+Compression involves efficiently extracting essential information from large datasets while removing redundant data, highlighting a model’s ability to understand the data’s underlying structure (Wei et al., 2024).
+LLMs are expected to perform this compression during training (Zhao et al., 2023).
+Initially, after random initialization, the data representations are chaotic, but as training progresses, they become organized, allowing the model to filter out unnecessary information.
+Thus, assessing an LLM’s compression capacity is vital for understanding its learning efficiency and representational power, which are crucial for practical applications and real-world deployment.
 […]
-To validate the effectiveness of the Matrix Nuclear-Norm, we first conducted preliminary experiments on two language models of differing sizes.
-The results indicated a consistent decrease in Matrix Nuclear-Norm values as model size increased, signifying enhanced compression capabilities.
-Subsequently, we performed inference experiments on two widely used benchmark datasets, AlpacaEval (Dubois et al., 2024) and Chatbot Arena (Chiang et al., 2024), which cover a diverse range of language generation tasks.
-These benchmarks facilitate a comprehensive assessment of model inference performance.
-Our experimental findings confirm that the Matrix Nuclear-Norm accurately measures model compression capabilities and effectively ranks models based on performance, demonstrating its reliability and efficiency in practical applications.
+To validate the Matrix Nuclear-Norm, we conducted preliminary experiments on two language models of different sizes.
+Results showed a consistent decrease in Matrix Nuclear-Norm values as model size increased, indicating enhanced compression capabilities.
+We also performed inference experiments on benchmark datasets, AlpacaEval (Dubois et al., 2024) and Chatbot Arena (Chiang et al., 2024), covering diverse language generation tasks.
+These benchmarks provide a comprehensive assessment of model inference performance.
+Our findings confirm that the Matrix Nuclear-Norm accurately measures model compression capabilities and ranks models based on performance, demonstrating its reliability and efficiency.
 Our empirical investigations yield the following insights:
-  1. 1.
-    Proposal of the Matrix Nuclear-Norm: We present a new method that leverages the nuclear norm, successfully reducing the computational complexity associated with evaluating language models from $O(n^{3})$ to $O(n^{2})$.
-This reduction minimizes dependence on SVD, making the Matrix Nuclear-Norm a more efficient alternative to Matrix Entropy.
-  2. 2.
-    Extensive Experimental Validation: We validated the effectiveness of the Matrix Nuclear-Norm on language models of various sizes.
-Results indicate that this metric accurately assesses model compression capabilities, with values decreasing as model size increases, reflecting its robust evaluation capability.
-  3. 3.
-    Benchmark Testing and Ranking: We conducted inference tests on widely used benchmark datasets, AlpacaEval and Chatbot Arena, evaluating the inference performance of models across different sizes and ranking them based on the Matrix Nuclear-Norm.
-The results demonstrate that this metric can efficiently and accurately evaluate the inference performance of medium and small-scale models, highlighting its broad application potential in model performance assessment.
+  • Proposal of the Matrix Nuclear-Norm: We introduce a method leveraging the nuclear norm, reducing computational complexity from $O(n^{3})$ to $O(n^{2})$.
+This reduction minimizes SVD dependence, making Matrix Nuclear-Norm a more efficient alternative to Matrix Entropy.
+  • Extensive Experimental Validation: We validated the Matrix Nuclear-Norm on language models of various sizes.
+Results show this metric accurately assesses model compression capabilities, with values decreasing as model size increases, reflecting its robust evaluation capability.
+  • Benchmark Testing and Ranking: We conducted inference tests on benchmark datasets, AlpacaEval and Chatbot Arena, evaluating inference performance across different model sizes and ranking them based on the Matrix Nuclear-Norm.
+Results demonstrate this metric efficiently and accurately evaluates medium and small-scale models, highlighting its broad application potential in model performance assessment.
 2 Related Work
-Evaluating Large Language Models: Metrics and Innovations.
-The evaluation of large language models (LLMs) is rapidly evolving, driven by diverse tasks, datasets, and benchmarks (Wang et al., 2023; Celikyilmaz et al., 2020; Zheng et al., 2023; Chang et al., 2024).
-Effective evaluation is crucial for enhancing model performance and reliability.
-Traditional metrics, such as accuracy, F1 score (Sasaki, 2007), BLEU (Papineni et al., 2002), and ROUGE (Lin, 2004), assess model outputs against annotated labels, while perplexity and cross-entropy loss focus on input data.
-Recent innovations, like Matrix Entropy (Wei et al., 2024), provide a means to evaluate information compression through activation matrix entropy, shedding light on internal representations.
-However, the computational intensity of Matrix Entropy, reliant on SVD, highlights the urgent requirement for developing more efficient evaluation methods in the context of LLMs.
-Scaling Laws and Performance Metrics.
-Recent studies have consistently shown that as models and datasets expand, performance metrics such as loss improve predictably (Hestness et al., 2017; Kaplan et al., 2020; Clark et al., 2022).
-This trend is observed across various domains, including NLP and computer vision (Hoffmann et al., 2022; Bahri et al., 2024).
-Consequently, both industry and academia have prioritized the development of larger models (Wei et al., 2022; Touvron et al., 2023).
-Notably, scaling laws reveal a power-law relationship in performance enhancement.
-In addition to traditional metrics like loss and task accuracy (Zhang et al., 2022; Dey et al., 2023; Biderman et al., 2023), we introduce the Matrix Nuclear-Norm as a novel metric for assessing information compression and redundancy reduction in LLMs.
-Insights from Information Theory for Large Language Models.
-Information theory has significantly enhanced our understanding of neural networks.
-The information bottleneck concept has been pivotal in elucidating supervised learning dynamics (Tishby et al., 2000).
-Recently, researchers have applied information-theoretic principles to self-supervised learning in vision tasks (Tan et al., 2023; Skean et al., 2023).
-Additionally, lossless compression techniques for LLMs, utilizing arithmetic coding, have been explored (Chen et al., 2024a; Valmeekam et al., 2023; Delétang et al., 2023).
-In this paper, we build on these foundations to introduce the Matrix Nuclear-Norm, an information-theoretic metric aimed at effectively evaluating the information compression and redundancy reduction capabilities of LLMs.
+LLM Evaluation and Scaling Laws.
+Evaluating large language models (LLMs) is a multifaceted challenge, as it requires capturing both task-specific performance and internal representational efficiency.
+Scaling laws have become a foundational framework for studying how LLM performance evolves with model size and data volume (Kaplan et al., 2020; Ruan et al., 2024).
+These studies demonstrate that model performance on tasks like language modeling and fine-tuning often follows predictable power-law relationships with respect to model parameters and dataset size, emphasizing the importance of scaling for achieving state-of-the-art results.However, scaling laws typically focus on external metrics such as cross-entropy loss, offering limited insight into how LLMs manage internal knowledge representation.
+For instance, the ability of LLMs to compress knowledge, eliminate redundancy, and retain structured information remains poorly understood with traditional methods.
+Addressing these gaps requires structural metrics that go beyond task outcomes to directly evaluate the internal embeddings and activation patterns of LLMs.
+LLM Evaluation Metrics.
+Traditional evaluation metrics such as perplexity, BLEU (Papineni et al., 2002), and ROUGE (Lin, 2004) primarily measure task-specific outcomes, assessing how well model outputs align with ground truth data.
+While these metrics are effective for evaluating surface-level outputs, they do not capture the underlying mechanisms of LLMs, such as the diversity or compression of embeddings.
+Similarly, accuracy and F1 score (Sasaki, 2007) focus on classification performance, making them less applicable to the generative tasks typical of LLMs.To bridge this gap, structural metrics such as Matrix Entropy have been introduced.
+Matrix Entropy (Wei et al., 2024) employs information theory to assess the entropy of covariance matrices derived from LLM embeddings.
+This metric evaluates how effectively a model removes redundancy and encodes structured information, offering a measure of its compression capabilities.
+For instance, Matrix Entropy can reveal differences in embedding distributions across models of varying sizes, reflecting their capacity to extract meaningful patterns from large datasets.
+However, its reliance on Singular Value Decomposition (SVD) results in a computational complexity of $O(n^{3})$, limiting its applicability to modern large-scale models.
+To overcome these limitations, we propose the Matrix Nuclear-Norm as a scalable alternative.
+By leveraging the $L_{1,2}$ norm as a convex approximation of matrix rank, the Matrix Nuclear-Norm reduces computational complexity to $O(n^{2})$.
+This makes it feasible for evaluating embeddings from large-scale LLMs while preserving the insights provided by Matrix Entropy, such as compression efficiency.
 3 Preliminaries
-This section presents the fundamental concepts and metrics used in our study to assess model performance, specifically focusing on discriminability, diversity, and the nuclear norm.
+This section presents the fundamental concepts for model performance evaluation: discriminability, diversity, and nuclear norm.
 3.1 Discriminability Measurement: F-NORM
-Higher discriminability corresponds to lower prediction uncertainty in the response matrix $A$, quantified using Shannon Entropy (Shannon, 1948):
-$H(A)=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{C}A_{i,j}\log(A_{i,j}).$ (1)
-In this equation, $B$ represents the number of samples, while $C$ denotes the number of distinct categories.
-Minimizing $H(A)$ results in maximal discriminability (Vu et al., 2019; Long et al., 2016), achieved when each row contains only one entry equal to 1 and all other entries equal to 0, indicating complete prediction certainty.
-Alternatively, we can focus on maximizing the Frobenius norm $\|A\|_{F}$:
+Higher discriminability corresponds to lower prediction uncertainty in the response matrix $A$.
+When $A$ is normalized as a probability matrix (i.e., $\sum_{j=1}^{C}A_{i,j}=1,\ \forall i\in[B]$), this uncertainty can be quantified using Shannon Entropy (Shannon, 1948):
+$H(A)=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{C}A_{i,j}\log\left(A_{i,j}\right)$ (1)
+where $B$ is the number of samples, $C$ the feature dimension, and $A_{i,j}$ the normalized activation value. Lower entropy indicates higher discriminability.
+An alternative measurement is the Frobenius norm:
 $\|A\|_{F}=\sqrt{\sum_{i=1}^{B}\sum_{j=1}^{C}|A_{i,j}|^{2}}.$ (2)
-The Frobenius norm serves as a measure of the overall magnitude of the entries in $A$ and also reflects prediction certainty.
-Theorem 1. Given a matrix $A$, $H(A)$ and $\|A\|_{F}$ are strictly inversely monotonic.
-The proof is detailed in Sect A.4 of the Appendix. Thus, minimizing $H(A)$ is equivalent to maximizing $\|A\|_{F}$. The bounds for $\|A\|_{F}$ are:
-$\sqrt{\frac{B}{C}}\leq\|A\|_{F}\leq\sqrt{B},$ (3)
-where the upper and lower bounds correspond to maximum and minimum discriminability, respectively.
+This norm reflects activation intensity, with higher values indicating more concentrated distributions.
+Theorem 1. For a row-normalized matrix $A\in\mathbb{R}_{+}^{B\times C}$ (i.e., $\sum_{j=1}^{C}A_{i,j}=1,\ \forall i$), $H(A)$ and $\|A\|_{F}$ are strictly inversely monotonic.
+The norm satisfies dimensional bounds:
+$\sqrt{\frac{B}{C}}\leq\|A\|_{F}\leq\sqrt{B}$ (3)
+where the lower bound achieves when $A$ has uniform distributions (maximal uncertainty), and the upper bound when $A$ contains one-hot vectors (minimal uncertainty).
+The proof is given in Appendix A.5.
 3.2 Diversity Measurement: Matrix Rank
-Prediction diversity involves the number of unique categories in matrix $A$, denoted as $C_{p}(A)$. The expected value $E_{C}$ represents the average number of categories:
-$E_{C}=E_{A\sim\mathcal{D}}\left(C_{p}(A)\right).$ (4)
-In practice, $C_{p}(A)$ can be approximated by the rank of $A$ when $\|A\|_{F}$ is near its upper bound:
-$C_{p}(A)\approx\operatorname{rank}(A).$ (5)
+In LLMs, diversity reflects the model’s ability to utilize its latent representation space effectively. For a given dataset $\mathcal{D}$, the expected diversity of outputs is defined as:
+$E_{C}=\mathbb{E}_{A\sim\mathcal{D}}\big{[}C_{p}(A)\big{]}$ (4)
+To approximate $C_{p}(A)$, we construct a sparse matrix $M\in\{0,1\}^{B\times C}$ where each row contains a one-hot vector indicating the argmax position:
+$M_{i,j}=\begin{cases}1,&j=\arg\max_{k}A_{i,k}\\ 0,&\text{otherwise}\end{cases}$ (5)
+The capacity measure then becomes:
+$C_{p}(A)=\mathrm{rank}\big{(}M\odot A\big{)}\approx\mathrm{rank}(A)$ (6)
+where $\odot$ denotes element-wise product.
+The maximum value of $C_{p}(A)$ is $\min(B,C)$, where $C$ is the output representation dimension.
+Maximizing $C_{p}(A)$ ensures effective utilization of the representation space, promoting robustness through reduced redundancy.
 3.3 Nuclear Norm
 The nuclear norm is an important measure related to diversity and discriminability.
-Theorem 2. When $\|A\|_{F}\leq 1$, the convex envelope of $\operatorname{rank}(A)$ is the nuclear norm $\|A\|_{\star}$. The theorem is proved in Fazel (2002).
-The nuclear norm $\|A\|_{\star}$, defined as the sum of singular values of $A$, has significant implications for assessing model performance. With $\|A\|_{F}\leq\sqrt{B}$, we have:
-$\frac{1}{\sqrt{D}}\|A\|_{\star}\leq\|A\|_{F}\leq\|A\|_{\star}\leq\sqrt{D}\cdot\|A\|_{F},$ (6)
-where $D=\min(B,C)$. Therefore, maximizing $\|A\|_{\star}$ ensures high diversity and discriminability.
-The upper bound of $\|A\|_{\star}$ is given by:
+Theorem 2. When $\|A\|\leq 1$ (where $\|A\|$ is the spectral norm), the convex envelope of $\operatorname{rank}(A)$ is the nuclear norm $\|A\|_{\star}$. The theorem is proved in Fazel (2002).
+For a matrix $A\in\mathbb{R}^{B\times C}$ with $\|A\|_{F}\leq\sqrt{B}$, let $D=\min(B,C)$. The relationships between the nuclear norm and Frobenius norm are:
+$\|A\|_{F}\leq\|A\|_{\star}\leq\sqrt{D}\cdot\|A\|_{F}.$ (7)
+Therefore, maximizing $\|A\|_{\star}$ encourages higher rank, which implies high diversity and discriminability. The upper bound of $\|A\|_{\star}$ is further bounded by:
 $\|A\|_{\star}\leq\sqrt{D\cdot B}.$ (8)
@@
 4.2 Matrix Nuclear-Norm
-Calculating the nuclear norm of matrix $A$ requires calculating its SVD, with a time complexity of $O\left(\min(B^{2}C,BC^{2})\right)$, simplifying to $O(n^{3})$, where $n$ represents the maximum of dimensions $B$ and $C$.
-While this computation is manageable for smaller dimensions, it becomes increasingly time-consuming as dimensions grow.
-Furthermore, SVD may not converge in some cases, it is thus important to develop more efficient approximations of singular values.
-Since $A$ is typically sparse, with only a few non-zero responses in each category, its singular values can be approximated by aggregating these non-zero responses.
-Theorem 3. When $\|A\|_{F}$ approaches the upper bound $\sqrt{B}$, the $j$-th largest singular value $\sigma_{j}$ can be approximated as:
-$\sigma_{j}\approx\text{top}\left(\sum_{i=1}^{B}A_{i,j}^{2},j\right),\quad\forall j\in\{1,\ldots,D\}.$ (8)
-The proof is provided in Sect. A.5 of the Supplementary Materials. Using this approximation, the batch nuclear norm can be efficiently calculated as:
-$\|\hat{A}\|_{*}=\sum_{j=1}^{D}\operatorname{top}\left(\sqrt{\sum_{i=1}^{B}A_{i,j}^{2}},j\right).$ (9)
+For a matrix $A\in\mathbb{R}^{B\times C}$, computing its exact nuclear norm via Singular Value Decomposition (SVD) requires $O(\min(B^{2}C,BC^{2}))$ time, which is equivalent to $O(n^{3})$ with $n=\max(B,C)$.
+While feasible for small matrices, this becomes computationally prohibitive for large-scale models.
+Additionally, numerical instability may arise in SVD computations for ill-conditioned matrices.
+Sparsity Prior: When $A$ exhibits column-wise sparsity (i.e., non-zero activations concentrate in a subset of columns), we can approximate its singular values by leveraging column norms.
+Let $\|A\|_{F}$ denote the Frobenius norm, bounded by $\|A\|_{F}\leq\sqrt{\min(B,C)}\cdot\sigma_{\max}(A)$, where $\sigma_{\max}(A)$ is the largest singular value.
+Theorem 3. (Column-Norm Approximation)
+If $A$ has rapidly decaying column norms $\{\|A_{:,j}\|_{2}\}_{j=1}^{C}$, the $j$-th largest singular value $\sigma_{j}(A)$ can be approximated by the $j$-th largest column norm:
+$\sigma_{j}(A)\approx\mathrm{Sort}\left(\left\{\|A_{:,j}\|_{2}\right\}_{j=1}^{C}\right)_{[j]},\quad j\in\{1,\dots,r\},$ (9)
+where $r=\mathrm{rank}(A)$. The proof is given in Sect. A.6 (Supplementary Materials). The nuclear norm is then approximated as:
+$\|\hat{A}\|_{\star}\approx\sum_{j=1}^{D}\mathrm{Sort}\left(\left\{\|A_{:,j}\|_{2}\right\}_{j=1}^{C}\right)_{[j]},$ (10)
+where $D\leq r$ is a hyperparameter controlling approximation precision, and $\widetilde{A}$ denotes the column-sparse approximation of $A$.
+Remark: This approximation holds under the assumption that off-diagonal correlations between columns are negligible (i.e., $A^{\top}A\approx\mathrm{diag}(\|A_{:,1}\|_{2}^{2},\dots,\|A_{:,C}\|_{2}^{2})$).
+For correlated columns, a diagonal correction term may be required.
 This approach indicates that the primary components of the $L_{1,2}$-norm can effectively approximate the nuclear norm when $\|A\|_{F}$ is close to $\sqrt{B}$, while other components can be considered noise.
-Compared to traditional SVD-based methods, such as Guo et al.
-(2015), our Matrix Nuclear-Norm calculation offers two main advantages: (1) it reduces computational complexity from $O(n^{3})$ to $O(n^{2})$, and (2) it relies solely on ordinary floating-point calculations, avoiding risks of non-convergence.
-The complete algorithm for Matrix Nuclear-Norm is detailed in Algorithm 1.
+Compared to traditional SVD-based methods (e.g., Guo et al.
+(2015)), this approach reduces computational complexity from $O(n^{3})$ to $O(n^{2})$ and avoids convergence issues by using only standard floating-point operations.
+The complete algorithm is detailed in Algorithm 1.
 Definition of Matrix Nuclear-Norm. The approach can ultimately be expressed as:
@@
 Algorithm 1 Algorithm of Matrix Nuclear-Norm
-0:  Sentence representations (hidden states from LLM) $\mathcal{S}=\left\{X_{i}\right\}_{i=1}^{m}$, where $x_{i}\in\mathbb{R}^{d\times 1}$, $d$ is the hidden dimension of representation, and $L_{\text{input}}$ is the length of the sentence.
-1:  $\mu=\frac{1}{m}\sum_{i=1}^{m}x_{i}$ // Calculate the mean embedding
-2:  $\mathbf{X}_{\text{norm}}=\frac{\mathbf{X_{i}}-\mu}{\|\mathbf{X_{i}}-\mu\|_{2,\text{row}}}$ // Normalize the activation matrix
+0:  Sentence representations $\mathcal{S}=\left\{X_{i}\right\}_{i=1}^{m}$, where $X_{i}\in\mathbb{R}^{d\times 1}$, $d$ is the hidden dimension, and $L_{\text{input}}$ is the sentence length.
+1:  $\mu=\frac{1}{m}\sum_{i=1}^{m}X_{i}$ // Mean embedding
+2:  $\mathbf{X}_{\text{norm}}=\frac{\mathbf{X}-\mu}{\|\mathbf{X}-\mu\|_{2,\text{row}}}$ // Normalize matrix
 […]
-[…]
+4:  $\Sigma_{D}=\{\sigma_{1},\sigma_{2},\dots,\sigma_{D}\}$ // Top $D$ norms
 […]
@@
 5 Experiments of Large Language Models
-5.1 Implementation Details
-5.1.1 Baselines
-Cross-Entropy Loss. For a text sequence $U=\{u_{1},\ldots,u_{T}\}$, the cross-entropy loss is defined as follows (Yang, 2019):
-$L(U)=-\frac{1}{T}\sum_{i=1}^{T}\log P(u_{i}|u_{<i};\Theta)$ (11)
-This loss quantifies the model’s ability to predict $u_{i}$ based on preceding words $u_{<i}$.
-We evaluate the performance of the Matrix Entropy metric in comparison to this baseline, utilizing the same datasets and models as referenced in (Kaplan et al., 2020).
-Perplexity. The perplexity of a language model for a text sequence $U=\{u_{1},\ldots,u_{T}\}$ is defined as (Neubig, 2017):
-$PPL(U)=\exp\left(-\frac{1}{T}\sum_{i=1}^{T}\log P(u_{i}|u_{<i};\Theta)\right)$ (12)
-Perplexity measures the model’s capacity to predict the next word, reflecting the average number of attempts required for correct predictions.
+The models and datasets used in this paper are thoroughly introduced in  A.2.
+5.1 Baselines
+Cross-Entropy Loss. Cross-entropy is a key metric for evaluating LLMs by measuring the divergence between predicted and true probability distributions. The formula is given as (Wei et al., 2024):
+$\mathcal{L}_{\text{CE}}=-\frac{1}{T}\sum_{i=1}^{T}\log P(u_{i}\mid u_{<i};\Theta)$ (12)
+where $u_{i}$ is the target token at position $i$, $P(u_{i}\mid u_{<i};\Theta)$ is the conditional probability predicted by the model, and $T$ is the sequence length.
+Lower values indicate better prediction accuracy.
+We compare this baseline with the Matrix Nuclear Norm metric, using the same datasets and models from (Kaplan et al., 2020).
+Perplexity. Perplexity measures how well a language model predicts a sequence of words. For a text sequence $\mathbf{U}=\{u_{1},\ldots,u_{T}\}$, it is defined as (Neubig, 2017; Wei et al., 2024):
+$\text{PPL}(\mathbf{U})=\exp\left(-\frac{1}{T}\sum_{i=1}^{T}\log P(u_{i}\mid u_{<i};\Theta)\right)$ (13)
+Lower perplexity indicates better performance, showing that fewer attempts are needed to predict the next token.
 Matrix Entropy of a Dataset.
-Given a dataset $\mathcal{D}=\{\mathcal{S}_{i}\}_{i=1}^{n}$, where $\mathcal{S}_{i}$ represents the samples (sentence embeddings) within the dataset, the (normalized) matrix entropy of $\mathcal{D}$ is defined as (Wei et al., 2024):
-$H(\mathcal{D})=\frac{\sum_{i=1}^{n}H(\Sigma_{\mathcal{S}_{i}})}{n\log d}.$ (13)
+For a dataset $\mathcal{D}=\{\mathbf{S}_{i}\}_{i=1}^{n}$, where $\mathbf{S}_{i}\in\mathbb{R}^{d\times d}$ represents sentence embedding covariance matrices, the normalized matrix entropy is defined as (Wei et al., 2024):
+$H(\mathcal{D})=\frac{1}{n\log d}\sum_{i=1}^{n}H\left(\frac{\sigma(\mathbf{S}_{i})}{\|\sigma(\mathbf{S}_{i})\|_{1}}\right)$ (14)
+where $\sigma(\mathbf{S}_{i})$ denotes the singular values of matrix $\mathbf{S}_{i}$, and $H(\cdot)$ is the Shannon entropy computed over the normalized singular value distribution.
 5.1.1 Language Models
@@
 The inclusion of these diverse models enriches our research perspective and facilitates an in-depth exploration of the inference performance and scaling laws of LLMs across different parameter sizes.
-5.1.3 Language Datasets
-Pretraining Datasets.
-All sizes of the Cerebras-GPT models were pretrained on the Pile dataset.
-Due to resource constraints, we randomly selected a subset of 10,000 data samples from Wikipedia (Foundation, 2024) and openwebtext2 (Skylion007, 2019) to observe the Matrix Nuclear-Norm, as these datasets are part of the Pile.
-In addition to the datasets used for model pretraining, we incorporated supplementary datasets that were not directly involved in the training process.
-Instruction-Tuning Datasets.
-To assess the effectiveness of the Matrix Nuclear-Norm, we selected the dolly-15k dataset (Conover et al., 2023), generated by human workers, as one of our instruction datasets.
-We specifically focused on the "context" portion of this dataset, which contains more informative text.
-This dataset will also be referenced in Section 5.3.2 for analyzing context length, prompt learning, and contextual operations.
-RLHF Dataset.
-We utilized the hh-rlhf dataset (Bai et al., 2022), which contains human preference data related to usefulness and harmlessness, as our Reinforcement Learning with Human Feedback (RLHF) dataset.
-Each entry includes a pair of texts, one marked as "chosen" and the other as "rejected." We input the "chosen" portions of the dataset into the model and subsequently calculated the performance of the Matrix Nuclear-Norm.
-Benchmark Datasets.
-To evaluate the performance of the Matrix Nuclear-Norm, we selected three benchmark datasets: OpenBookQA (Mihaylov et al., 2018), Winogrande (Sakaguchi et al., 2021), and PIQA (Bisk et al., 2020).
-These benchmarks are structured in a multiple-choice format.
-We used each question along with its correct answer as input to compute the respective evaluation metrics.
-Prompt Learning Datasets.
-We conducted prompt learning experiments using the OpenOrca dataset (Lian et al., 2023b), where each question-answer (QA) pair features a unique system prompt.
-To explore the impact of these prompts on Matrix Nuclear-Norm, we carefully selected three generally instructive prompts (see Table 13) to observe their roles in the QA pairs.
-Validation Datasets.
-Additionally, we performed inference and evaluation on two widely used datasets: AlpacaEval (Dubois et al., 2024) and Chatbot Arena (Chiang et al., 2024).
-We conducted inference on both datasets, evaluated the responses using our proposed metrics, and ranked the models accordingly.
 5.2 Matrix Nuclear-Norm Observation
-5.2.1 A Comparative Analysis of Computational Time
+5.2.1 Comparing Computational Time
 Figure 1: Cerebras-GPT: Time comparison
@@
 For example, the  [^1] M model showed that Matrix Nuclear-Norm was 8.58 times quicker than Matrix Entropy.
-The significant efficiency gain is attributed to the lower computational complexity of Matrix Nuclear-Norm, $O(m\cdot n+n\log n)$, compared to Matrix Entropy’s $O(n^{3})$.
-This makes Matrix Nuclear-Norm not only an effective but also a highly efficient metric for evaluating LLMs, especially as models continue to scale in size.
+The significant efficiency gain is due to the lower complexity of Matrix Nuclear-Norm, $O(m\cdot n+n\log n)$, versus Matrix Entropy’s $O(n^{3})$, where $m$ is the embedding dimension (columns).
+This makes it an efficient metric for LLM evaluation, especially for large-scale models.
 In summary, Matrix Nuclear-Norm achieves comparable evaluation accuracy to Matrix Entropy but with vastly superior computational efficiency, making it a practical and scalable choice for assessing LLMs.
-Model Size Matrix Entropy Time (s) Matrix Nuclear-Norm Time (s) Ratio
- [^1] M 623.5367 72.6734 8.5800
-256M 1213.0604 110.8692 10.9414
-590M 2959.6949 184.7785 16.0175
-1.3B 6760.1893 379.0093 17.8365
-2.7B 12083.7105 732.6385 16.4934
-6.7B 38791.2035 1598.4151 24.2685
-13B 59028.4483 2984.1529 19.7806
-Table 1: CEREBRAS-GPT: Time Comparison between Matrix Entropy and Matrix Nuclear-Norm
+Model Size ME Time(s) MNN Time(s) Ratio
+ [^1] M 623.5 72.7 8.6
+256M 1213.1 110.8 10.9
+590M 2959.7 184.8 16.0
+1.3B 6760.2 379.0 17.8
+2.7B 12083.7 732.6 16.5
+6.7B 38791.2 1598.4 24.3
+13B 59028.4 2984.2 19.8
+Table 1: Cerebras-GPT: Time Comparison between Matrix Entropy (ME) and Matrix Nuclear-Norm (MNN)
 5.2.2 Scaling Law of Matrix Nuclear-Norm
 To affirm Matrix Nuclear-Norm’s efficacy as an evaluative metric, we evaluated Cerebras-GPT models on four datasets including dolly-15k, Wikipedia, openwebtext2, and hh-rlhf comparing Matrix Nuclear-Norm, matrix entropy, perplexity, and loss.
-Results, detailed in Table 11 (Appendix), demonstrate Matrix Nuclear-Norm’s consistent decrease with model size enlargement, signifying better data compression and information processing in larger models.
-This trend (see in Figure 2(d)) validates Matrix Nuclear-Norm’s utility across the evaluated datasets.
+As shown in Table 10, Matrix Nuclear-Norm decreases consistently with model size, indicating better data compression and processing in larger models.
+This trend (Figure 2(b)) validates Matrix Nuclear-Norm’s utility across datasets.
 Notably, anomalies at the 2.7B and 13B highlight areas needing further exploration.
-(a) Cross-Entropy Loss
-(b) Perplexity
-(c) Matrix Entropy
-(d) Matrix Nuclear-Norm
-Figure 2: Comparison of Matrix Nuclear-Norm, matrix entropy, loss, and perplexity when model scales up.
+(a) Matrix Entropy
+(b) Matrix Nuclear-Norm
+Figure 2: Comparison of Matrix Nuclear-Norm, matrix entropy when model scales up.
 5.2.3 Relationship of Benchmark Indicators
@@
 5.3.1 Sentence Operation Experiments
-Figure 3 clearly indicates that sentence manipulations significantly influence Matrix Nuclear-Norm values, which generally decline as model size increases.
-This trend confirms the enhanced information compression capabilities of larger models.
-The ranking of Matrix Nuclear-Norm values by operation is as follows: Reverse $>$ Shuffle & Reverse $>$ Shuffle $>$ Base.
-This indicates that disrupting sentence structure through Reverse and Shuffle & Reverse operations leads to higher Matrix Nuclear-Norm values due to increased information chaos and processing complexity.
-In contrast, the Shuffle operation has minimal effect on compression, while the Base condition consistently yields the lowest Matrix Nuclear-Norm values, signifying optimal information compression efficiency with unaltered sentences.
-Despite the overall downward trend in Matrix Nuclear-Norm values with increasing model size, the 2.7B model exhibits slightly higher values for Shuffle and Base operations compared to the 1.3B model.
-This anomaly suggests that the 2.7B model may retain more nuanced information when handling shuffled data or operate through more intricate mechanisms.
-However, this does not detract from the overarching conclusion that larger models excel at compressing information, thereby demonstrating superior processing capabilities.
+Figure 3 shows sentence manipulations impact Matrix Nuclear-Norm values.
+These values decrease with model size, in line with established scaling laws similar to those governing matrix entropy and perplexity (PPL).
+As models grow larger, they can capture data patterns more efficiently, reducing redundant information representation, which directly lowers the nuclear norm.
+The ranking Reverse $>$ Shuffle & Reverse $>$ Shuffle $>$ Base reflects input disruption.
+Reverse flips the sentence, introducing maximum disorder and causing a large norm increase.
+Shuffle only partially rearranges elements, leading to a smaller rise.
+The unaltered Base condition enables optimal compression.
+Notably, the 2.7B model has slightly higher Shuffle and Base values than the 1.3B model, yet this doesn’t challenge the conclusion that larger models compress better.
+The norm increases with text length because longer texts carry more information, increasing entropy and computational complexity.
+More data means more potential redundancy for the model to process, driving up the norm value.
+These results clarify model behavior in relation to size, input structure, and length.
 Figure 3: Results of sentence operation. Shuffling and reversing disrupt the text structure and diminish the informational content, leading to an increase in Matrix Nuclear-Norm.
 5.3.2 Analysis of Length Dynamics
-Figure 4: The Matrix Nuclear-Norm values for contexts of varying lengths show that as text length increases, the Matrix Nuclear-Norm continues to rise and tends to converge.
-The analysis reveals that Matrix Nuclear-Norm values generally increase as input length rises, aligning with our expectations (see Figure 4).
-Longer inputs necessitate that the model manage and compress more information, which naturally leads to higher Matrix Nuclear-Norm values.
+Figure 4: The Matrix Nuclear-Norm values increase consistently with longer text input lengths, reflecting the model’s ability to capture more information.
+The analysis reveals that Matrix Nuclear-Norm generally increase as input length rises, aligning with our expectations (see Figure 4).
+Longer inputs necessitate that the model manage and compress more information, which naturally leads to higher Matrix Nuclear-Norm.
 Most models exhibit this trend, indicating effective handling of the increased information load.
-However, the gpt-2.7B and gpt-13B models display anomalies in their Matrix Nuclear-Norm values at 64 and 128 tokens, where the value at 128 tokens is lower than that at 64 tokens.
+However, the Cerebras-GPT-2.7B and Cerebras-GPT-13B models display anomalies in their Matrix Nuclear-Norm values at 64 and 128 tokens, where the value at 128 tokens is lower than that at 64 tokens.
 This discrepancy may be attributed to these models employing different information compression mechanisms or optimization strategies tailored to specific input lengths, allowing for more effective compression at those lengths.
-Overall, aside from a few outliers, the results largely conform to expectations, demonstrating that Matrix Nuclear-Norm values increase with input length, reflecting the greater volume and complexity of information that models must handle.
-To address the observed trend of rising Matrix Nuclear-Norm values with longer sentences, we incorporated a normalization step in our methodology via dividing the Matrix Nuclear-Norm values by the sentence length.
+Overall, aside from a few outliers, the results largely conform to expectations, demonstrating that Matrix Nuclear-Norm values increase with input length, reflecting the greater volume and complexity of information that models must handle.To address the observed trend of rising Matrix Nuclear-Norm values with longer sentences, we incorporated a normalization step in our methodology via dividing the Matrix Nuclear-Norm values by the sentence length.
 This adjustment helps mitigate any biases introduced by models that tend to generate longer sentences during inference.
-GPT MODEL SIZE
-LENGTH  [^1] M 256M 590M 1.3B 2.7B 6.7B 13B
-64 0.4574 0.4125 0.3787 0.3486 0.4053 0.3315 0.4148
-128 0.5293 0.4680 0.4270 0.3835 0.4143 0.3477 0.4032
-512 0.7883 0.6978 0.6251 0.5554 0.5265 0.4468 0.4422
-1024 0.9132 0.8787 0.7802 0.6953 0.6351 0.5383 0.5028
-Table 2: Analysis of Length Dynamics
 5.3.3 Analysis of Prompt Learning
-The experimental results (shown in Table 3) indicate that we performed inference on different sizes of GPT models using three carefully selected prompts (shown in Table 13) and calculated the Matrix Nuclear-Norm values of their responses.
+The experimental results (shown in Table 2) indicate that we performed inference on different sizes of Cerebras-GPT models using three carefully selected prompts (shown in Table 12) and calculated the Matrix Nuclear-Norm values of their responses.
 As the model size increased, the Matrix Nuclear-Norm values gradually decreased, demonstrating that larger models possess greater information compression capabilities.
 The prompts significantly influenced Matrix Nuclear-Norm, with variations reflecting the models’ responses to prompt complexity.
-Specifically, GPT-1.3B showed a notable decrease in Matrix Nuclear-Norm after the input prompts, indicating its sensitivity to them, while GPT-2.7B exhibited smaller changes.
-In contrast, GPT-6.7B displayed minimal variation across all prompts, suggesting stable performance regardless of prompt detail.
+Specifically, Cerebras-GPT-1.3B showed a notable decrease in Matrix Nuclear-Norm after the input prompts, indicating its sensitivity to them, while Cerebras-GPT-2.7B exhibited smaller changes.
+In contrast, Cerebras-GPT-6.7B displayed minimal variation across all prompts, suggesting stable performance regardless of prompt detail.
 Overall, more detailed prompts resulted in larger information volumes in the model’s responses, leading to corresponding changes in Matrix Nuclear-Norm values.
-Table 3: Results of prompt learning with (Empty Prompt) and without (Prompt 1, 2, 3) the use of prompts.
+Table 2: Results of prompt learning without Prompt and with (Prompt 1, 2, 3) the use of prompts.
 Incorporating prompts as prefixes before the QA pairs enhances the models’ ability to achieve better compression.
@@
 Cerebras-GPT-6.7B 0.132042 0.128346 0.124094 0.133211 0.12923 ↓0.002812
-6 Implementing Proposed Metrics: Evaluating and Ranking Language Models in Practice
+6 Evaluating and Ranking LLMs
 6.1 Inference-Based Model Assessment
@@
 The analysis revealed that Matrix Nuclear-Norm reliably ranks model performance, with lower values indicating enhanced information processing efficiency, particularly as model size scales up.
-For instance, the Llama-3 70B model demonstrated superior compression capabilities compared to its 8B counterpart, as reflected by significantly lower Matrix Nuclear-Norm values across both benchmarks (see Table A.2 in the Appendix).
+For instance, the Llama-3 70B model demonstrated superior compression capabilities compared to its 8B counterpart, as reflected by significantly lower Matrix Nuclear-Norm values across both benchmarks (see Table 7).
 A similar trend was observed in the Vicuna family, where Matrix Nuclear-Norm values consistently decreased from 0.4623 for the 7B model to 0.3643 for the 33B model on the AlpacaEval dataset, indicating progressive improvements in information handling (see Table 3).
@@
 Overall, these results substantiate Matrix Nuclear-Norm as a robust and reliable tool for evaluating and ranking LLMs, demonstrating its capacity to capture critical aspects of model performance across diverse benchmarks.
-Model DataSet 7B 13B 33B
+Model Data 1.3B 6.7B 7B
+DeepSeek Alpaca 0.4882 0.3472 0.3352
+Arena 0.5754 0.4175 0.4357
 Vicuna Alpaca 0.4623 0.4159 0.3643
 Arena 0.4824 0.4311 0.3734
-Model DataSet 1.3B 6.7B 7B
-DeepSeek Alpaca 0.4882 0.3472 0.3352
-Arena 0.5754 0.4175 0.4357
 Table 3: Matrix Nuclear-Norms in Vicuna and DeepSeek Responses
-6.2 Matrix Nuclear-Norm Benchmarking: Ranking Mid-Sized Models
+6.2 Matrix Nuclear-Norm for Model Ranking
 In this experimental section, we utilized Matrix Nuclear-Norm to evaluate the responses of LLMs, focusing on 7B and 70B variants.
@@
 These results highlight that factors such as architecture, training methodology, and data complexity play crucial roles in information processing capabilities beyond scale.
-MODEL Matrix Nuclear-Norm Rank
+     MODEL Matrix Nuclear-Norm
 Alpaca Arena-Hard Avg Score
-DeepSeek-7B 0.3352 0.4357 0.3855 $\downarrow$
-Gemma-7B 0.3759 0.3998 0.3879 $\downarrow$
-Vicuna-7B 0.4623 0.4824 0.4724 $\downarrow$
-LLaMA 2-7B 0.4648 0.5038 0.4843 $\downarrow$
-QWEN 1.5-7B 0.4866 0.5165 0.5016 $\downarrow$
-Mistral-7B 0.4980 0.5126 0.5053 $\downarrow$
-QWEN 2-7B 0.5989 0.5751 0.5870 $\downarrow$
-QWEN 1.5-72B 0.5291 0.5065 0.5178 $\downarrow$
-QWEN 2-72B 0.5261 0.4689 0.4975 $\downarrow$
-Llama 3-70B 0.4935 0.4967 0.4951 $\downarrow$
-Llama 2-70B 0.3862 0.4086 0.3974 $\downarrow$
-Table 5: Matrix Nuclear-Norm Rankings: A Comparative Analysis of Model Performance
+QWEN 2-7B 0.5989 0.5751 0.5870
+Mistral-7B 0.4980 0.5126 0.5053
+QWEN 1.5-7B 0.4866 0.5165 0.5016
+LLaMA 2-7B 0.4648 0.5038 0.4843
+Vicuna-7B 0.4623 0.4824 0.4724
+Gemma-7B 0.3759 0.3998 0.3879
+DeepSeek-7B 0.3352 0.4357 0.3855
+QWEN 1.5-72B 0.5291 0.5065 0.5178
+QWEN 2-72B 0.5261 0.4689 0.4975
+Llama 3-70B 0.4935 0.4967 0.4951
+Llama 2-70B 0.3862 0.4086 0.3974
+Table 4: Descending Competence Rankings via Matrix Nuclear Norm: Small and Large LMs
 To validate the design rationale and robustness of the Matrix Nuclear-Norm, we conducted a series of ablation studies.
@@
 This not only confirms its applicability across diverse models but also verifies its stability and reliability in handling large-scale datasets.
-We also provide an ablation study in the appendix, further proving the method’s efficiency and accuracy in evaluating LLMs.
+We also provide an ablation Cerebras-GPT: study in the appendix, further proving the method’s efficiency and accuracy in evaluating LLMs.
 7 Conclusion
@@
 8 Limitations
-Although Matrix Nuclear-Norm performs well in evaluating the performance of LLMs, it still has some limitations.
-First, since Matrix Nuclear-Norm’s computation relies on the model’s hidden states, the evaluation results are sensitive to both the model architecture and the training process.
-As a result, under different model designs or training settings, especially for models like GPT-1.3B and GPT-2.7B, inconsistencies in Matrix Nuclear-Norm’s performance may arise, limiting its applicability across a wider range of models.
-Additionally, while Matrix Nuclear-Norm offers computational efficiency advantages over traditional methods, it may still face challenges with resource consumption when evaluating extremely large models.
-As model sizes continue to grow, further optimization of Matrix Nuclear-Norm’s computational efficiency and evaluation stability is required.
+Although Matrix Nuclear-Norm (MNN) performs well in evaluating LLM performance, it has three main limitations.
+First, as MNN computation relies on hidden states, the results are sensitive to model architecture and training processes.
+This may cause performance inconsistencies across different model designs or training settings (particularly between Cerebras-GPT-1.3B and Cerebras-GPT-2.7B), potentially limiting broader applicability.
+Second, while MNN offers computational advantages over traditional methods, it may still face resource challenges when evaluating extremely large models, requiring further optimization for scalability.
+Third, our current implementation uses MNN primarily as an evaluation metric rather than a training objective.
+However, we recognize its potential for analyzing information compression dynamics during training, which could provide valuable insights into model optimization.
+Future work should explore this direction while addressing the method’s sensitivity to architectural variations.
+Notably, despite observed anomalies in specific configurations, MNN demonstrates consistent computational efficiency and accuracy across various model sizes and data sampling strategies.
+We will enhance our discussion of these performance variations to better clarify the method’s robustness boundaries and operational constraints.
+These limitations highlight the need for continued research into architecture-agnostic evaluation frameworks and optimized computation strategies as language models scale.
 9 Ethics Statement
@@
 This commitment reflects our dedication to responsible AI research and the broader implications of using such data in language model development.
+10 Reproducibility
+We emphasize the importance of reproducibility in the development and evaluation of our newly proposed metric, Matrix Nuclear-Norm.
+To facilitate reproducibility, we provide detailed information regarding our data processing and parameter settings:
+Data Processing and Parameter Settings: We outline the preprocessing steps applied to each dataset, ensuring that other researchers can accurately replicate our methodology.
+All hyperparameters and configuration settings used during the experiments are specified in the code, offering clarity on the experimental conditions.
+Experimental Procedures: We detail the specific steps required to evaluate the Matrix Nuclear-Norm, including its application to each dataset and the metrics used for performance assessment.
+Code Availability: Our implementation code, evaluation scripts, and pretrained models will be made publicly available upon acceptance of this paper, enabling others to reproduce our experiments and validate our findings.
+By adhering to these guidelines, we aim to ensure that our work is accessible and reproducible for future research endeavors.
 Appendix A Appendix
@@
 A.1.1 Different Model Family
+(a) Cross-Entropy Loss
+(b) Perplexity
+Figure 5: Comparison of loss, and perplexity when model scales up.
 In addition to evaluating Matrix Nuclear-Norm within the Cerebras-GPT model series, we extended our experiments to the Pythia model family, which spans from 14M to 12B parameters and is trained on consistent public datasets.
@@
 These findings confirm the Matrix Nuclear-Norm as a reliable metric for large-scale datasets, effectively evaluating information compression and redundancy elimination in LLMs.
-Table 6: Ablation study of differnet sampling strategies on the Wikimedia[18] dataset.
+Table 5: Ablation study of different sampling strategies on the Wikimedia(Foundation, 2024) dataset.
 MODEL SAMPLING STRATEGY STANDARD DEVIATION
@@
 CEREBRAS-GPT-1.3B 0.5684 0.5670 0.5676 0.5699 0.5693 0.0004975
-A.2 Supplementary Experiment Results
+A.2 Model Selection and Datasets for Analysis
+Model Selection.
+To investigate language model scaling, we employed a diverse set of transformer-based large language models (LLMs) across varying parameter sizes.
+A key focus of our analysis was the Cerebras-GPT model (Gao et al., 2020), which ranges from 111 million to 13 billion parameters, providing a comprehensive look at scaling effects in pre-trained models.
+Additionally, we included scaled versions of the Pythia model (Biderman et al., 2023), with parameter counts ranging from 14 million to 12 billion, enabling a broader analysis of model performance across different scales.
+To ensure a well-rounded evaluation, we also tested a variety of models, including the DeepSeek series (1.3B, 6.7B, 7B) (Guo et al., 2024), Llama3 series (8B, 70B) (Dubey et al., 2024), QWEN 2 series (0.5B, 1.5B, 7B, 72B) (Yang et al., 2024), and Vicuna models (7B, 13B, 33B) (Chiang et al., 2023).
+For additional comparative insights, we included models of similar scale, such as Gemma-7B (Team et al., 2024) and Mistral-7B (Jiang et al., 2023).
+Datasets for Analysis.
+Our experiments were conducted using several key benchmark datasets.
+We selected AlpacaEval(Dubois et al., 2024) and ChatBot Arena (Zheng et al., 2023) as the primary datasets for model evaluation.
+Additionally, subsets from Wikipedia (Foundation, 2024) and OpenWebText2 (Skylion007, 2019) were utilized to track variations in Matrix Nuclear-Norm values, especially with the Cerebras-GPT models.
+To validate the Matrix Nuclear-Norm metric, we employed the dolly-15k dataset (Conover et al., 2023) for instruction tuning and the hh-rlhf dataset (Bai et al., 2022) for reinforcement learning with human feedback (RLHF).
+Further evaluations were performed on benchmark datasets such as OpenBookQA (Mihaylov et al., 2018), Winogrande (Sakaguchi et al., 2021), and PIQA (Bisk et al., 2020).
+Lastly, prompt learning experiments with the OpenOrca dataset (Lian et al., 2023b) provided a comprehensive framework for assessing the Matrix Nuclear-Norm’s effectiveness across a variety of inference tasks.
+A.3 Supplementary Experiment Results
 The following results provide additional insights into the Matrix Nuclear-Norm evaluations and comparisons across various language models:
   1. 1.
-    Tables A.2 and A.2 present the Matrix Nuclear-Norm evaluation results during the inference process for Llama-3 and QWEN-2.
+    Tables 7 and 6 present the Matrix Nuclear-Norm evaluation results during the inference process for Llama-3 and QWEN-2.
   2. 2.
@@
   3. 3.
-    Table 11 contains the complete results for the comparison of Matrix Nuclear-Norm and other metrics based on Cerebras-GPT family considered in Figure 2(d).
+    Table 10 contains the complete results for the comparison of Matrix Nuclear-Norm and other metrics based on Cerebras-GPT family considered in Figure 2(b).
   4. 4.
@@
     Table 12 shows the prompts used for the investigation of prompt learning.
-Figure 6: Pythia: Time Comparison of Matrix Entropy and Nuclear-Norm
-Model Size Matrix Entropy Time (s) Matrix Nuclear-Norm Time (s) Ratio
-14M 52.8669 22.2652 2.3772
-31M 114.0820 28.1842 4.0477
-70M 320.6641 24.3188 13.1855
-160M 631.9762 41.6187 15.1817
-410M 1040.9764 80.9814 12.8481
-1B 4650.1264 114.0639 40.8387
-1.4B 6387.0392 347.8670 18.3858
-2.8B 8127.1343 343.3888 23.6778
-6.9B 28197.8172 816.6332 34.5350
-12B 47273.5235 1276.1128 37.0485
-Table 7: Pythia Model: Matrix Entropy vs. Matrix Nuclear-Norm Time Comparison
 Model DataSet 0.5B 1.5B 7B 72B
@@
 Arena 0.5817 0.4967
-Table 9: Matrix Nuclear-Norm in Llama3 esponses
+Table 7: Matrix Nuclear-Norm in Llama 3 Responses
+Figure 7: Pythia: Time Comparison of Matrix Entropy and Nuclear-Norm
+Model Size ME (s) MNN (s) Ratio
+14M 52.9 22.3 2.4
+31M 114.1 28.2 4.0
+70M 320.7 24.3 13.2
+160M 632.0 41.6 15.2
+410M 1040.9 81.0 12.8
+1B 4650.1 114.1 40.8
+1.4B 6387.0 347.9 18.4
+2.8B 8127.1 343.4 23.7
+6.9B 28197.8 816.6 34.5
+12B 47273.5 1276.1 37.0
+Table 8: Pythia Model: Matrix Entropy(ME) vs. Matrix Nuclear-Norm(MNN) Time Comparison
 GPT MODEL SIZE
@@
 PIQA MATRIX NUCLEAR-NORM 0.4868 0.4327 0.4164 0.3826 0.4452 0.3675 0.4149
-Table 10: Language modeling indicators on openbookqa, winogrande and piqa.
+Table 9: Language modeling indicators on openbookqa, winogrande and piqa.Except for the matrix nuclear norm, the data is sourced from Wei et al. (2024)
+Table 10: The table illustrates the performance metrics for a range of GPT models on the Dolly-15k, Wikipedia, OpenWebText2, and HH-RLHF datasets, encompassing matrix entropy, loss, and perplexity.
+Except for the matrix nuclear norm, the data is sourced from Wei et al.
+(2024), underscoring the relationship between model scale and its performance.
 DATASET INDICATORS GPT MODELS SIZE
@@
 MATRIX NUCLEAR-NORM 0.6309 0.5716 0.5307 0.4771 0.4959 0.4277 0.4518
-Table 11: Language modeling indicators on the Dolly-15k, Wikipedia, OpenWebText2, and HH-RLHF datasets.
-Table 12: The language modeling indicators (where lower values indicate better performance) for the Pythia models were evaluated on the Dolly-15k, Wikipedia, OpenWebText2, and HH-RLHF datasets.
+Table 11: Language modeling indicators for Pythia models across Dolly-15k, Wikipedia, OpenWebText2, and HH-RLHF datasets (lower values indicate better performance).
+Except for the matrix nuclear norm, data is derived from Wei et al.
+(2024), showcasing the correlation between model scale and performance.
 PYTHIA MODELS SIZE
@@
 HH-RLHF MATRIX NUCLEAR-NORM 0.7353 0.7674 0.6964 0.6182 0.5886 0.4825 0.5141 0.4839 0.4562 0.4481
-Table 13: The prompts selected from OpenOrca[30] dataset.
 Prompt ID Prompt Content
@@
 Prompt 3 You are an AI assistant. User will give you a task. Your goal is to complete the task as faithfully as you can. While performing the task think step-by-step and justify your steps.
-A.3 Analysis of Algorithmic Complexity
+Table 12: The prompts selected from OpenOrca(Lian et al., 2023b) dataset.
+GPT MODEL SIZE
+LENGTH  [^1] M 256M 590M 1.3B 2.7B 6.7B 13B
+64 0.4574 0.4125 0.3787 0.3486 0.4053 0.3315 0.4148
+128 0.5293 0.4680 0.4270 0.3835 0.4143 0.3477 0.4032
+512 0.7883 0.6978 0.6251 0.5554 0.5265 0.4468 0.4422
+1024 0.9132 0.8787 0.7802 0.6953 0.6351 0.5383 0.5028
+Table 13: Analysis of Length Dynamics
+A.4 Analysis of Algorithmic Complexity
 The primary computational expense of Matrix Nuclear-Norm arises from the calculation and sorting of the L2 norm of the matrix.
@@
 We will support this claim with detailed complexity analysis and experimental results.
-A.3.1 Time Complexity Analysis
+A.4.1 Time Complexity Analysis
 Analysis 1: Time Complexity of Matrix Entropy
@@
 This indicates that Matrix Nuclear-Norm is computationally more efficient, especially as $n$ increases.
-A.3.2 Experimental Validation and Comparative Analysis
+A.4.2 Experimental Validation and Comparative Analysis
 To empirically validate the theoretical time complexities, we conducted experiments using matrices of various sizes.
@@
 Impact of Constant Factors Although both $O(n^{2})$ and $O(n^{3})$ indicate asymptotic behavior, Matrix Nuclear-Norm’s significantly smaller constant factors make it computationally favorable even for moderately sized matrices, as evidenced in our experimental results.
-A.3.3 Conclusion of the Complexity Analysis
+A.4.3 Conclusion of the Complexity Analysis
 Through this detailed analysis and experimental validation, we conclude the following:
@@
   • Experimental results confirm that Matrix Nuclear-Norm offers superior time efficiency for evaluating large-scale models, particularly those with millions or billions of parameters.
-A.4 Proof of Theorem 1: Detailed Version
-Here we present the proof of Theorem 1, demonstrating the strictly inverse monotonic relationship between $H(A)$ and $\|A\|_{F}$.
-Given a matrix $A\in\mathbb{R}^{B\times C}$, where each element is non-negative and each row sums to 1, i.e.,
-$\sum_{j=1}^{C}A_{i,j}=1\quad\text{for}\ i=1,2,\ldots,B.$
-We define the Shannon entropy $H(A)$ as:
-$H(A)=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{C}A_{i,j}\log(A_{i,j}),$
-and the Frobenius norm $\|A\|_{F}$ as:
-$\|A\|_{F}=\sqrt{\sum_{i=1}^{B}\sum_{j=1}^{C}|A_{i,j}|^{2}}.$
-Step 1: Analysis for a Fixed Row Vector
-For a fixed row $i$, let $a_{j}=A_{i,j}$ such that $a_{j}\geq 0$ and $\sum_{j=1}^{C}a_{j}=1$. The entropy for this single row is given by:
-$H_{i}=-\sum_{j=1}^{C}a_{j}\log a_{j},$
-and the Frobenius norm for this row vector $a$ is:
-$\|a\|_{2}=\sqrt{\sum_{j=1}^{C}a_{j}^{2}}.$
-To determine when $H_{i}$ is maximized or minimized, we apply the method of Lagrange multipliers. Consider the function:
-$L(a_{1},a_{2},\ldots,a_{C},\lambda)=-\sum_{j=1}^{C}a_{j}\log a_{j}+\lambda\left(\sum_{j=1}^{C}a_{j}-1\right).$
-Taking partial derivatives with respect to $a_{j}$ and setting them to zero:
+A.5 Proof of Theorem 1
+We prove the strictly inverse monotonic relationship between the entropy $H(A)$ and the Frobenius norm $\|A\|_{F}$ for a non-negative matrix $A\in\mathbb{R}^{B\times C}$ where each row represents a probability distribution:
+$\sum_{j=1}^{C}A_{i,j}=1,\quad A_{i,j}\geq 0,\quad\forall i=1,\ldots,B.$
+Definitions:
+  • Entropy:
+    $H(A)=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{C}A_{i,j}\log(A_{i,j})$
+  • Frobenius norm:
+    $\|A\|_{F}=\sqrt{\sum_{i=1}^{B}\sum_{j=1}^{C}A_{i,j}^{2}}$
+Step 1: Single-Row Analysis
+For a row $\mathbf{a}=[a_{1},\ldots,a_{C}]$ with $\sum_{j}a_{j}=1$:
+  • Row entropy: $H_{i}=-\sum_{j=1}^{C}a_{j}\log a_{j}$
+  • Row norm: $\|\mathbf{a}\|_{2}=\sqrt{\sum_{j=1}^{C}a_{j}^{2}}$
+Extrema via Lagrange Multipliers:
+The Lagrangian $L=-\sum_{j}a_{j}\log a_{j}+\lambda(\sum_{j}a_{j}-1)$ yields:
 $\frac{\partial L}{\partial a_{j}}=-\log a_{j}-1+\lambda=0\implies a_{j}=e^{\lambda-1}.$
-Since $\sum_{j=1}^{C}a_{j}=1$, we find $a_{j}=\frac{1}{C}$ for all $j$. Thus, $H_{i}$ is maximized when all $a_{j}$ are equal. In this case,
-$\|a\|_{2}=\sqrt{\sum_{j=1}^{C}\left(\frac{1}{C}\right)^{2}}=\sqrt{\frac{1}{C}}.$
-Conversely, $H_{i}$ reaches its minimum when one element $a_{k}=1$ and all other $a_{j}=0$. At this point,
-$\|a\|_{2}=\sqrt{1^{2}+0}=1.$
-This demonstrates that $H_{i}$ and $\|a\|_{2}$ are inversely related for each row.
-Step 2: Inverse Monotonicity for the Entire Matrix
-Since $H(A)$ is the average of all $H_{i}$ values:
-$H(A)=\frac{1}{B}\sum_{i=1}^{B}H_{i},$
-and $\|A\|_{F}$ is the square root of the sum of the squared norms $\|a\|_{2}^{2}$ for each row:
-$\|A\|_{F}=\sqrt{\sum_{i=1}^{B}\|a\|_{2}^{2}},$
-the inverse relationship between $H_{i}$ and $\|a\|_{2}$ implies that $H(A)$ and $\|A\|_{F}$ are also inversely related across the entire matrix $A$.
-Step 3: Determining the Bounds of $\|A\|_{F}$
-To find the bounds, first note that when $a_{j}=\frac{1}{C}$ for all $j$,
-$\|A\|_{F}=\sqrt{\sum_{i=1}^{B}\sum_{j=1}^{C}\left(\frac{1}{C}\right)^{2}}=\sqrt{B\cdot\frac{1}{C}}=\sqrt{\frac{B}{C}},$
-which corresponds to the minimum discriminability.
-For the upper bound, when each row has only one element equal to 1,
-$\|A\|_{F}=\sqrt{\sum_{i=1}^{B}1}=\sqrt{B},$
-indicating maximum discriminability.
-Step 4: Practical Implications
-The inverse monotonic relationship between $H(A)$ and $\|A\|_{F}$ is significant in machine learning and classification tasks, where maximizing $\|A\|_{F}$ implies achieving greater certainty in predictions.
-This property can be used to evaluate model performance, where higher $\|A\|_{F}$ corresponds to a model that is more confident and discriminative in its predictions.
-Therefore, $H(A)$ and $\|A\|_{F}$ are strictly inversely monotonic, and the proof is complete.
-A.5 Proof of Theorem 3
-Given that $\|A\|_{F}$ is near $\sqrt{B}$, we want to approximate the $j$-th largest singular value $\sigma_{j}$ as $\text{top}\left(\sum_{i=1}^{B}A_{i,j}^{2},j\right)$.
-This approximation will be derived by examining the contributions of the columns of $A$.
+Normalization gives $a_{j}=\frac{1}{C}$, achieving:
+  • Maximum entropy: $H_{i}=\log C$
+  • Minimum norm: $\|\mathbf{a}\|_{2}=\sqrt{\frac{1}{C}}$
+Minimum entropy occurs when $a_{k}=1$ (one-hot vector):
+  • Minimum entropy: $H_{i}=0$
+  • Maximum norm: $\|\mathbf{a}\|_{2}=1$
+Monotonicity: For fixed $C$, $H_{i}$ and $\|\mathbf{a}\|_{2}$ are strictly inversely monotonic (shown via derivative analysis or majorization theory).
+Step 2: Matrix-Level Generalization
+For the full matrix:
+  • $H(A)=\frac{1}{B}\sum_{i=1}^{B}H_{i}$
+  • $\|A\|_{F}=\sqrt{\sum_{i=1}^{B}\|\mathbf{a}_{i}\|_{2}^{2}}$
+Key Observation: If each row’s entropy $H_{i}$ decreases (increases), its norm $\|\mathbf{a}_{i}\|_{2}$ increases (decreases).
+Thus: - $\|A\|_{F}^{2}=\sum_{i=1}^{B}\|\mathbf{a}_{i}\|_{2}^{2}$ decreases (increases) as $H(A)$ increases (decreases).
+Step 3: Norm Bounds
+Maximum $\|A\|_{F}$: When all rows are one-hot:
+$\|A\|_{F}=\sqrt{B}$
+Minimum $\|A\|_{F}$: When all rows are uniform:
+$\|A\|_{F}=\sqrt{\frac{B}{C}}$
+Step 4: Implications for LLMs
+The inverse monotonicity implies:
+  • High $\|A\|_{F}$: Concentrated predictions (low entropy, high confidence).
+  • Low $\|A\|_{F}$: Dispersed predictions (high entropy, high diversity).
+Thus, $\|A\|_{F}$ serves as a proxy for evaluating LLM confidence-diversity tradeoffs.
+Conclusion
+The strict inverse monotonicity between $H(A)$ and $\|A\|_{F}$ is rigorously established, justifying $\|A\|_{F}$ as a metric for LLM evaluation.
+A.6 Proof of Theorem 3
+Assuming $\|A\|_{F}\approx\sqrt{B}$ and the columns of $A$ are approximately orthogonal, we approximate the $j$-th largest singular value $\sigma_{j}$ as the $j$-th largest column norm of $A$.
+Formally,
+$\sigma_{j}\approx\text{top}\left(\sqrt{\sum_{i=1}^{B}A_{i,j}^{2}},\ j\right),$
+where $\text{top}(S,j)$ denotes the $j$-th largest element in set $S$. This approximation holds under the following analysis:
 1.
-Decomposition of $A$ and Eigenvalues: By singular value decomposition (SVD), $A=U\Sigma V^{T}$, where $\Sigma$ is a diagonal matrix containing singular values $\sigma_{1},\sigma_{2},\ldots,\sigma_{D}$.
-When looking at the product $A^{T}A$, the element at position $(j,j)$ is:
-$(A^{T}A)_{j,j}=\sum_{i=1}^{B}A_{i,j}^{2}$
-This sum represents the total squared magnitude of the entries in the $j$-th column of $A$.
+Decomposition and Gram Matrix: Let $A=U\Sigma V^{T}$ be the SVD of $A$, where $\Sigma=\text{diag}(\sigma_{1},\dots,\sigma_{D})$ with $D=\min(B,C)$.
+The diagonal entries of the Gram matrix $A^{T}A$ are:
+$(A^{T}A)_{j,j}=\sum_{i=1}^{B}A_{i,j}^{2}=\|\mathbf{a}_{j}\|_{2}^{2},$
+where $\mathbf{a}_{j}$ is the $j$-th column of $A$.
 2.
-Connecting to Singular Values: Since singular values represent the contribution of orthogonal projections of $A$, if $\|A\|_{F}$ is near $\sqrt{B}$, the main contribution to the nuclear norm $\|A\|_{*}$ comes from columns with the largest sums of squared values $\sum_{i=1}^{B}A_{i,j}^{2}$.
-3.
-Approximation: For matrices with well-distributed entries in each column (a common property when $\|A\|_{F}\approx\sqrt{B}$), the top singular values can be effectively estimated by sorting these column norms.
-Hence, the largest singular values $\sigma_{j}$ approximately correspond to the largest column norms $\sqrt{\sum_{i=1}^{B}A_{i,j}^{2}}$.
-Therefore, for each $j\in\{1,\ldots,D\}$,
-$\sigma_{j}\approx\text{top}\left(\sum_{i=1}^{B}A_{i,j}^{2},j\right).$
-This concludes the proof that the $j$-th largest singular value $\sigma_{j}$ can be approximated by the top $j$ values of column-wise norms of $A$, as stated in Theorem 3.
+Relating Column Norms to Singular Values: When columns of $A$ are nearly orthogonal, $\sigma_{j}\approx\|\mathbf{a}_{j}\|_{2}$.
+Under $\|A\|_{F}\approx\sqrt{B}$, the nuclear norm $\|A\|_{\star}=\sum_{j=1}^{D}\sigma_{j}$ is dominated by the largest column norms.
+3. Singular Value Approximation: For matrices with low column-wise correlations, the $j$-th singular value satisfies:
+$\sigma_{j}\approx\text{top}\left(\{\|\mathbf{a}_{k}\|_{2}\mid 1\leq k\leq C\},\ j\right).$
+4. Efficient Nuclear Norm Approximation: The batch nuclear norm is approximated as:
+$\|\hat{A}\|_{\star}=\sum_{j=1}^{D}\text{top}\left(\{\|\mathbf{a}_{k}\|_{2}\},\ j\right).$
+This approximation is valid when $A$ has approximately orthogonal columns, a condition implied by $\|A\|_{F}\approx\sqrt{B}$.